### v4

# Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior

Abstract
Power system operators and electric utility companies often charge consumers a peak demand charge when the aggregate system demand reaches its maximum, a practice known as coincident peak shaving. These charges incentivize consumers to reduce usage during critical periods, alleviating stress on electricity transmission and distribution systems, while also helping to recover the grid investment costs. In this paper, we analyze the problem through the lens of game theory, developing a model that captures how strategic consumer behavior influences overall system efficiency. Our results reveal that the coincident peak shaving game can be concave, quasi-concave with discontinuities, or non-concave with discontinuities, depending on the extent of consumers’ demand-shifting capabilities. In a two-agent, two-period framework, we derive closed-form Nash equilibrium solutions for each scenario and extend our analysis to multi-agent settings. We prove the stability of these equilibrium points and propose an algorithm to compute equilibrium outcomes across all game configurations. Furthermore, we show that while the decentralized game model achieves peak-shaving performance comparable to a centralized approach, it does so at the cost of increased anarchy. In scenarios characterized by quasi-concave and non-concave conditions, our analytical results demonstrate that anarchy grows with greater consumer flexibility and disparities in marginal shifting costs, and we explore how the number of agents affects system efficiency. Finally, numerical simulations are provided to validate our theoretical findings.
Index Terms:
Coincident peak shaving, Bottleneck game, Price of anarchy, Equilibrium stability, Demand response, Power system operation
I Introduction
Managing peak electricity demand is crucial in power system transmission and distribution due to the need to maintain an instantaneous balance between supply and demand. At the transmission level, total demand must not exceed the maximum generation capacity. At the distribution level, where a tree-branched network of transformer substations steps down voltages, the load at each tier must remain below the capacity of the respective transformers. When demand exceeds the system’s generation or substation capacity, operators may be forced to disconnect portions of the grid to cut off the connecting demand to ensure the secure operation of the remaining system, potentially leading to significant financial losses and adverse impacts on the life quality [Kirschen and Strbac, 2018].
With the advent of smart electric meters capable of recording time-dependent electricity usage and the widespread deployment of distributed energy resources (DER), power system operators and electric utility companies are implementing new pricing schemes to promote demand flexibility and alleviate peak demand stress. One such approach is time-of-use (ToU) tariffs [Salt River Project, 2025], in which utilities set hourly electricity prices to encourage consumers to shift consumption: increasing usage during low-priced periods (typically at night) and reducing it during high-priced periods (typically during the day). However, ToU tariffs mainly focus on recovering the cost of energy supply and shifting daily consumption, and they are less effective at curbing peak demand over longer periods, such as on a monthly basis, because the prices are not set high enough to further incentivize significant peak demand reduction [Trabish, 2016, Burger et al., 2020].
Utility companies may also impose direct peak demand charges, billing consumers based on their highest power usage with a prefixed peak price over an extended period—typically a month. These charges not only motivate consumers to reduce their peak consumption but also help recover distribution grid investment and maintenance costs, as the delivered grid capacity is determined by a consumer’s peak demand [Mullendore, 2017, Shi et al., 2017]. For instance, Consolidated Edison, serving New York City, may charge more than \$40 per kW of peak demand, with the peak being determined by the highest two consecutive 15-minute intervals in any monthly billing cycle. This effectively translates to a cost of approximately \$80 per kWh during peak periods, which is significantly higher than the standard energy supply rate of around 20 to 30 cents per kWh [Consolidated Edison, 2025a].
Yet, allocating peak demand charges becomes non-trivial when multiple consumers share a capacity bottleneck, such as a constrained substation, because the system’s peak does not directly correspond to each individual consumer’s peak [Chen et al., 2020b]. Charging a consumer for their peak demand may be inequitable if the overall system is not under peak stress, and such charges might not accurately represent the true system peak demand. This issue is particularly critical for utility companies, which are often publicly owned or regulated private monopolies with profit margins capped by their investment costs. Therefore, the revenue from peak demand charges should accurately reflect grid investment expenses based on the actual delivery capacity.
Coincident peak (CP) demand charges bill consumers based on their usage during the system’s peak period, offering an alternative that more accurately reflects the system condition and the corresponding investment costs. These charges are added to future electricity bills; for example, Texas’s 4CP mechanism applies charges based on a consumer’s average demand during the Electric Reliability Council of Texas (ERCOT) system’s peak hour in each of the four summer months (from June to September), with these charges affecting the following year’s bills. Notably, 4CP charges can represent up to 30% of an organization’s monthly electricity costs [Ögelman, 2016]. In 2023, the 4CP charge was approximately \$67/kWh [Grid Status, 2024], with the total coincident peak demand across the ERCOT region reaching 83.6 GW [ERCOT, 2025b], translating into an estimated \$5.6 billion payments. According to ERCOT’s annual demand response report, the 4CP mechanism facilitated a 3,500 MW demand reduction during peak periods, resulting in savings of approximately \$234.5 million, or 4.2% of the total 4CP-related costs [Potomac Economics, 2023]. A significant challenge, however, is that the peak period is identified only in hindsight. While numerous academic and industrial efforts have focused on accurately predicting the system’s peak period [Patrick and Wolak, 2001, Dowling et al., 2018, CPower, 2016], these studies often overlook the interactions among consumers, whose collective behavior determines the system peak. This gap naturally motivates a game-based framework and raises a critical question: Is a game-based framework workable for the CP shaving problem, and how does strategic consumer behavior affect system efficiency? Addressing this question is the main focus of this paper.
I-A Summary of main contributions and implications
In this work, we formulate a novel game framework for the CP shaving problem. As a first step toward applying a game-theoretic model to address CP shaving challenges, we consider a simplified two-agent, two-period setting, with extensions to more than two agents discussed in Section VI. The two agents represent clusters of customers or dominant oligopoly customers. The rationale for the two-period setting stems from its ability to represent a peak period and an off-peak period [Anunrojwong et al., 2024], as identified from historical market data, including demand shifting patterns and price signals. For example, power system peak demand typically occurs in the early evening, with off-peak periods covering the rest of the day. In the 4CP program in Texas, historical records in 2024 suggest that certain periods, such as around 17:00, are more likely to experience CP events [ERCOT, 2025b]. Our model focuses on minimizing system peak demand while deemphasizing off-peak performance, aligning with the primary objective of operators and utilities. This simplified framework offers both insight and tractability for analyzing the fundamental interactions between agents and their impact on system performance. Furthermore, as agents can access the system demand profile from publicly available market information [ERCOT, 2025a], they can infer the aggregated response of other agents. We assume that each agent observes the other’s aggregated demand distribution, allowing us to isolate strategic interactions from the complexities of predicting individual valuations. By balancing simplicity with realism, this approach lays a solid foundation for understanding CP shaving strategies and agent behaviors.
With our settings in place, we formulate the CP game model with a fixed CP charge price and an individual penalty term for demand shifting to determine the cost, where agents compete to determine the CP period and their corresponding demand strategies. The structure of the CP game model is influenced by the agents’ demand-shifting capabilities (flexibility), which are determined by their shifting costs and baseline demands. Agents with larger baseline demand differences between two periods and lower shifting costs exhibit greater flexibility. Depending on these factors, the model may exhibit concave, quasi-concave with discontinuities, or non-concave with discontinuities characteristics. By contrast, the conventional centralized peak shaving model is formulated as a straightforward convex optimization problem.
We then analytically derive the pure strategy Nash equilibrium (NE) for all CP game structures under the two-agent, two-period setting. Our analysis reveals that the NE is determined by the system’s unbalanced demand and the agents’ shifting capabilities. By treating the system as a switched dynamic system, we prove that the equilibrium of the game model is globally uniformly asymptotically stable, regardless of switching, provided all agents’ demand shifting is within the baseline limit. Furthermore, we demonstrate that a gradient-based algorithm with an update rule serving as a finite difference approximation of the asymptotically stable process can compute the equilibrium point.
We analyze the impact of gaming agents’ strategic behaviors by examining the peak shaving effectiveness and the price of anarchy (PoA). Our findings analytically indicate that the peak shaving effectiveness of the game model matches that of the centralized model but comes at the cost of economic inefficiency (anarchy), except for the concave game, where the outcomes are fully equivalent. For quasiconcave and non-concave games, we prove that the PoA increases with the disparities among agents, as measured by their marginal shifting costs. Additionally, under identical system conditions, we show that greater agent flexibility exacerbates PoA.
We also generalized our findings with more than two agents, we prove the unique NE under concave and quasiconcave conditions. For non-concave CP games, we show that system demand can still be balanced across two periods and derive the NE at the group level of agents. We also establish the stability of the equilibrium point and validate the effectiveness of the gradient-based algorithm. The NE reveals equivalent peak shaving effectiveness between the game and the centralized model while showing that changes in the number of agents influence the game type and, consequently, the PoA.
For industry practitioners and policymakers, our paper offers the following takeaways
  • We show that the game model with a stable NE is computationally efficient, and applying this model to the CP shaving problem does not compromise peak shaving performance. Thus, operators can adopt the game model without concerns about its impact on CP shaving effectiveness.
  • We identify that system’s inefficiencies in game environments stem from agents’ greater flexibility and disparities in their marginal shifting costs. This insight guides operators and policymakers in designing targeted incentive mechanisms tailored to agents, enhancing both system efficiency and equity.
  • We show that the economic performance of small systems with fewer agents gaming on CP shaving is sensitive to agents’ flexibility. Therefore, one should select flexible agents to form larger systems while leaving inflexible agents to form small gaming systems.
I-B Related works
To precisely position our study within the gap in the existing literature, we elaborate on our work from three perspectives closely related to our CP shaving problem: (i) Load management, (ii) CP demand charge, and (iii) Related game formulations.
Load management. Load management is a border concept covering peak demand reduction, load shifting, and demand response. Many studies, from the utility’s perspective, formulate optimization problems to design pricing mechanisms - such as tariffs or peak demand charges - aimed at minimizing convex costs or maximizing concave profits [Siano, 2014]. These pricing design approaches often rely on consumers’ price response behavior, typically leading to a two-layer framework (or Stackelberg game model): the upper layer determines the pricing strategy, while the lower layer models demand based on consumer response behavior [Chen et al., 2020a]. The lower layer may involve constructing utility functions to represent consumer preferences [Pennings and Smidts, 2003, Chen and Xu, 2024] or applying data-driven methods to learn behavior from historical price and consumption data [Jain et al., 2017]. From a demand-side perspective, the key challenge lies in distinguishing between baseline demand and observed demand to better understand consumers’ demand reduction decision-making behavior [Agrawal and Yücel, 2022]. Beyond individual consumers, the interactive influence of demand strategies has been widely studied through non-cooperative game formulations and generally achieves peak shaving [Joshi et al., 2021]. However, while these load management approaches contribute to peak demand reduction, they primarily focus on maximizing the profit of different market participants rather than directly addressing CP demand reduction in accordance with CP charging mechanisms.
CP demand charge. Academic studies and industrial solutions for CP shaving generally focus on predicting the timing of the system peak and making decisions from the consumers’ perspective. Prediction involves estimating the CP period as a probabilistic distribution, often using machine learning tools that leverage input features such as historical demand and weather conditions [Carmona et al., 2024, Dowling et al., 2018]. Following this, decisions are made based on the prediction results. Industrial solutions primarily emphasize short-term prediction, employing auto-regressive methods to iteratively update models and issue warning signals to consumers [CPower, 2016]. Academic solutions typically frame the problem as a scheduling task, making deliberate decisions based on CP period distributions. Examples include stochastic sequential optimization [Liu et al., 2013] and optimization with neural networks trained as decision policies [Dowling and Zhang, 2019]. Recently, an approach combining prediction and decision-making, termed ’decision-focused learning’ or contextual optimization [Sadana et al., 2024], has gained traction. This method trains weighting parameters using decision losses from downstream optimization rather than prediction losses, aligning with the optimization tasks to produce effective decisions [Elmachtoub and Grigas, 2022]. However, it has yet to be applied to the CP shaving problem. Moreover, existing studies often assume a single agent or ’price-taker’ condition, neglecting the impact of various agents’ decisions on CP period predictions. This motivates us to incorporate consumer interactions as a critical aspect of CP shaving design.
Related game formulations. Without specific constraints, the CP shaving problem can be abstracted as agents betting on discrete events, with winners sharing benefits based on their bids. This framework mirrors the classic concept of pari-mutuel betting, in which the odds (or CP periods) fluctuate in response to all participants’ betting strategies [Quandt, 1986, Willis, 1964]. Pari-mutuel betting is commonly applied in sports, where optimal betting strategies—both in size and target—are determined by solving discrete decision problems aimed at maximizing expected returns through the best combination of picks [Kenter, 2016, Rosner, 1975].
The CP shaving problem also shares similarities with congestion games by treating peak demand consumption as a form of resource usage, where increased utilization by multiple agents results in higher costs [Rosenthal, 1973]. Congestion games can often be formulated as potential games, guaranteeing the existence of pure-strategy Nash equilibria (NE) even in weighted settings, provided certain conditions on the utility functions are met [Harks and Klimm, 2016]. The primary difference between congestion games and CP shaving is in their cost structures: in congestion games, costs are determined by the aggregate resource usage, whereas in CP shaving, costs are driven by the peak-period demand—a single, critical resource. This makes the problem more analogous to a bottleneck congestion game, where the objective is to optimize the performance of the worst-performing element [Banner and Orda, 2007]. In such settings, pure-strategy NE has also been shown to exist, although they typically yield a non-optimal price of anarchy (PoA) [Busch and Magdon-Ismail, 2009].
Despite the valuable insights provided by these game-theoretic frameworks, the CP shaving problem exhibits key differences. Specifically, demand shifting can negatively impact consumer comfort and lead to profit losses, the cost structure is dictated by peak-period demand rather than the maximum cost across periods, and the overall system demand must remain unchanged. We address these gaps by integrating these considerations into our model and developing a CP shaving game framework.
II Model and Preliminaries
In this section, we formulate the CP shaving game model and introduce its definitions. We begin with a two-agent (x and y), two-period (1 and 2) model, where the two agents can represent clusters of customers with different baseline demand conditions, and the two periods correspond to a peak period and an off-peak period, respectively. We denote the agent x’s baseline demand in periods 1 and 2 as $X_{1}$ and $X_{2}$, respectively, with $x$ representing its demand shifting. Similarly, we denote the agent y’s baseline demand as $Y_{1}$ and $Y_{2}$ with $y$ as its demand shifting. The system’s baseline demand, given by summing both agents’ baseline demands, is denoted as $S_{\mathrm{b},1},S_{\mathrm{b},2}$, while the system’s total demand, which includes both agents’ demand shifting, is denoted as $S_{1},S_{2}$, For simplicity, we often omit the $x,y$ when there is no ambiguity. Under this two-period setting, we define the direction of demand shifting as follows: shifting demand to period 1 (away from period 2) is considered positive, while shifting demand to period 2 (away from period 1) is considered negative, e.g., for agent x is $X_{1}+x$ and $X_{2}-x$. Depending on the agents’ demand-shifting behaviors, the roles of the peak and off-peak periods can interchange, i.e., either period 1 or period 2 can become the peak or off-peak period. We use $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}\in\mathbb{R}^{+}$ to denote the penalty parameter for agent x and y, representing the comfort loss or perceived cost of shifting demand. These parameters can be identified from shifting behavior suggested by historical data. The fixed CP charge price is denoted by $\pi\in\mathbb{R}^{+}$ and $I(x,y)$ is the step function (indicator function). Both agent gaming to reduce CP demand while maximizing their payoff, and the game model is defined as $G=(N,\mathcal{X},U)$, where
  • $N=\{\mathrm{x},\mathrm{y}\}$ is the two-agent agent set;
  • $\mathcal{X}=\mathcal{X}_{\mathrm{x}}\times\mathcal{X}_{\mathrm{y}}$ is the strategy set formed by the product topology of each agent’s strategy set, where $\times$ is the topology product;
  • $U=\{f_{\mathrm{x}},f_{\mathrm{y}}\}$ is the payoff function set, where $f_{\mathrm{x}},f_{\mathrm{y}}:\mathcal{X}\rightarrow\mathbb{R}$.
During the game, agent x chooses strategy $x\in\mathcal{X}_{\mathrm{x}}$ to maximize its payoff $f_{\mathrm{x}}(x,y)$ with $y\in\mathcal{X}_{\mathrm{y}}$, and the payoff is to minimize its two periods’ costs,
$\max_{x}$ $f_{\mathrm{x}}(x,y)=-\pi(X_{1}+x)I(S_{1}(x,y)-S_{2}(x,y))-\pi(X_{2}-x)I(S_{2}(x,y)-S_{1}(x,y))-\alpha_{\mathrm{x}}x^{2},$ (1a)
$I(x)=\begin{cases}1&x\geq 0\\ 0&x<0\end{cases},$ (1b)
$S_{1}(x,y)=X_{1}+Y_{1}+x+y=S_{\mathrm{b},1}+x+y,$ (1c)
$S_{2}(x,y)=X_{2}+Y_{2}-x-y=S_{\mathrm{b},2}-x-y.$ (1d)
Agent y’s payoff function $f_{\mathrm{y}}(x,y)$ follows the same structure by changing baseline demand to $Y_{1},Y_{2}$, demand shifting to $y$, and penalty parameter to $\alpha_{\mathrm{y}}$. Note that there is always a CP period even if two periods of demand are equal; here, we treat period 1 as the CP period, reflected as the equal sign of the step function (1b). Also, we focus on demand shifting and assume a constant energy rate besides the peak demand charge, so we don’t consider the energy cost.
It is obvious that the game is compact, convex, and bounded, where compact means each $\mathcal{X}_{\mathrm{x}},\mathcal{X}_{\mathrm{y}}$ is compact, convex means each $\mathcal{X}_{\mathrm{x}},\mathcal{X}_{\mathrm{y}}$ is convex, and bounded means each $f_{\mathrm{x}},f_{\mathrm{y}}$ is bounded. Note that we use a quadratic penalty to express a soft constraint to make the strategy set compact, and we have the linear correlation of the number of variables and agents. We then introduce the definitions of equilibrium and continuity.
Definition 1.
We define the following: (1) Pure-strategy Nash equilibrium (Nash [Nash Jr, 1950]). $(x^{*},y^{*})\in\mathcal{X}$ is a NE in pure strategies of the game $G$ if and only if (iff) $f_{\mathrm{x}}(x^{*},y^{*})\geq f_{\mathrm{x}}(x,y^{*})$ and $f_{\mathrm{y}}(x^{*},y^{*})\geq f_{\mathrm{y}}(x^{*},y)$ for every $x\in\mathcal{X}_{\mathrm{x}},y\in\mathcal{X}_{\mathrm{y}}$.
(2) Upper semi-continuity (u.s.c.) and lower semi-continuity (l.s.c.). The function $f:\mathcal{X}\rightarrow\mathbb{R}$ is called upper semi-continuous (u.s.c.) or lower semi-continuous (l.s.c.) if for every $x_{0}$ such that
$\limsup\limits_{x\rightarrow x_{0}}f(x)\leq f(x_{0}),\text{ or }\liminf\limits_{x\rightarrow x_{0}}f(x)\geq f(x_{0}),$ (2)
for all $x$ in some neighborhood of $x_{0}$, respectively.
The step function makes the game analysis non-trivial. We thus separate the overall system into two periods corresponding to CP period in 1 and 2, i.e., $\mathcal{X}_{\mathrm{cp}1}=\{x,y|S_{1}(x,y)\geq S_{2}(x,y)\}$ and $\mathcal{X}_{\mathrm{cp}2}=\{x,y|S_{1}(x,y)<S_{2}(x,y)\}$, with the payoff functions, take agent x for example,
$f_{\mathrm{x},1}(x)=-\pi(X_{1}+x)-\alpha_{\mathrm{x}}x^{2},x\in\mathcal{X}_{\mathrm{cp}1},$
$f_{\mathrm{x},2}(x)=-\pi(X_{2}-x)-\alpha_{\mathrm{x}}x^{2},x\in\mathcal{X}_{\mathrm{cp}2},$ (3a)
It is obvious that each payoff function is concave, and by applying the first-order optimality condition [Boyd and Vandenberghe, 2004], we have
$x^{\prime}=\arg\max_{x}f_{\mathrm{x},1}(x)=-\frac{\pi}{2\alpha_{\mathrm{x}}},x\in\mathcal{X}_{\mathrm{cp}1},$
$x^{\prime}=\arg\max_{x}f_{\mathrm{x},2}(x)=\frac{\pi}{2\alpha_{\mathrm{x}}},x\in\mathcal{X}_{\mathrm{cp}2},$ (3b)
With these results, we define critical points, balance points, and system average demand in the game model.
Definition 2.
Critical points, balance points, and system average demand.
We define the following:
  1. 1.
    Critical point. $r_{\mathrm{x}}$ is the critical point of the payoff function $f_{\mathrm{x},2}$ described in (3), and the critical point of the payoff function $f_{\mathrm{x},1}$ is $-r_{\mathrm{x}}$. $r_{\mathrm{y}}$ follows the same idea.
    $r_{\mathrm{x}}=\frac{\pi}{2\alpha_{\mathrm{x}}},r_{\mathrm{y}}=\frac{\pi}{2\alpha_{\mathrm{y}}},$ (4a)
  2. 2.
    Agent balance point. $b_{\mathrm{x}},b_{\mathrm{y}}$ are the demand difference of agent x and y in two period periods
    $b_{\mathrm{x}}=\frac{X_{2}-X_{1}}{2},b_{\mathrm{y}}=\frac{Y_{2}-Y_{1}}{2},$ (4b)
  3. 3.
    System balance point. $b$ is the demand difference of the overall system in two period periods
    $b=b_{\mathrm{x}}+b_{\mathrm{y}}=\frac{S_{\mathrm{b},2}-S_{\mathrm{b},1}}{2}.$ (4c)
  4. 4.
    System average demand. $S$ is the average baseline demand of the overall system in two period periods is
    $S=\frac{S_{\mathrm{b},2}+S_{\mathrm{b},1}}{2}.$ (4d)
The critical point indicates the maximum demand each agent can shift to avoid the CP charge within one period. The agent balance allows each agent to shave its demand flat in two periods, which is the economic maximum demand shifting in the entire system. The reason is that once the agent balance point is reached, the CP charge is constant (half of the demand) regardless of the opponent’s strategy. Thus, $\min\{r_{\mathrm{x}},b_{\mathrm{x}}\}$ or $\max\{-r_{\mathrm{x}},b_{\mathrm{x}}\}$ define the maximum shifting capability of agent x. The system average demand is useful to show the CP charge when the system demand is balanced in the two period periods. We then define the capable and non-capable agents.
Definition 3.
Capable and non-capable agents. According to Definition 2, given agent x with baseline demand $X_{1},X_{2}$, CP charge $\pi$, and penalty parameter $\alpha_{\mathrm{x}}$, the agent x is capable if it satisfy
$-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},$ (5)
Define the agent that didn’t satisfy (5) as non-capable agent, including upper non-capable agent with $b_{\mathrm{x}}>r_{\mathrm{x}}$ and lower non-capable agent with $b_{\mathrm{x}}<-r_{\mathrm{x}}$.
According to Definition 2, (5) means agent x is economically capable of balancing its demand in the two periods during the game $G$ to reduce the CP charge. The reason is that the agent’s balance point $b_{\mathrm{x}}$ is within their critical point $[-r_{\mathrm{x}},r_{\mathrm{x}}]$ of both periods, indicating that they will reach the balance point before the critical point when shifting demand.
III Equilibrium analysis of the coincident peak game
In this section, we first show the properties of the two-agent two-period CP game $G$ with the specific agent type, and prove the pure-strategy NE of the CP game $G$.
III-A Game properties
We first show the property of CP game $G$. Determined by the parameters of all agents, the game performs differently.
Proposition 4.
Concave, quasiconcave/discontinuous, and non-concave/discontinuous CP game. Given critical point $r_{\mathrm{x}},r_{\mathrm{y}}$ and balance point $b_{\mathrm{x}},b_{\mathrm{y}},b$ as define in Definition 2, the two-agent two-period CP game $G$ satisfies one of the following:
  1. 1.
    Concave CP game. $G$ is concave that all agents’ payoff function $f_{\mathrm{x}}(x,y)$ ($f_{\mathrm{y}}(x,y)$) is concave in $x\in\mathcal{X}_{\mathrm{x}}$ ($y\in\mathcal{X}_{\mathrm{y}}$) for each $y\in\mathcal{X}_{\mathrm{y}}$ ($x\in\mathcal{X}_{\mathrm{x}}$) under the conditions of
    $b$ $<-r_{\mathrm{x}}-r_{\mathrm{y}},\text{if }S_{\mathrm{b},1}\geq S_{\mathrm{b},2},$ (6a)
    $b$ $>r_{\mathrm{x}}+r_{\mathrm{y}},\text{if }S_{\mathrm{b},1}<S_{\mathrm{b},2},$ (6b)
  2. 2.
    Quasiconcave CP game. $G$ is quasiconcave/discontinuous that all agents’ payoff function $f_{\mathrm{x}}(x,y)$ ($f_{\mathrm{y}}(x,y)$) is quasiconcave in $x\in\mathcal{X}_{\mathrm{x}}$ ($y\in\mathcal{X}_{\mathrm{y}}$) for each $y\in\mathcal{X}_{\mathrm{y}}$ ($x\in\mathcal{X}_{\mathrm{x}}$[Reny, 1999] under the conditions of
    $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},\ -r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}};$ (6c)
  3. 3.
    Non-concave CP game. $G$ is non-concave/discontinuous if it is not concave or quasiconcave. The condition is
    $\{-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}\}\cap$ [$[\{b_{\mathrm{x}}<-r_{\mathrm{x}}\}\cup\{b_{\mathrm{x}}>r_{\mathrm{x}}\}\cup\{b_{\mathrm{y}}<-r_{\mathrm{y}}\}\cup\{b_{\mathrm{y}}>r_{\mathrm{y}}\}].$ (6d)
Sketch of the proof.
The payoff function is quadratic if only one possible CP period during the gaming process, indicating a concave and continuous game. The indicator function discontinues the game when the CP period changes. According to the definition of the quasiconcave function, we separate three cases: 1) Agent x can switch the CP period before reaching their critical points in both periods 1 and 2; 2) Agent x can reach the critical point in period 2 before switching, and the function is l.s.c. as defined in Definition 1; 3) Agent x can reach the critical point in period 1 before switching, and the function is u.s.c. as defined in Definition 1. We show only the first case is able for both agents’ functions to be quasiconcave. Then the non-concave condition is derived from the complementary set of concave and quasiconcave conditions. Also, note that if one agent’s payoff function is quasiconcave or non-concave, the payoff functions of the other agents must also be quasiconcave or non-concave. The detailed proof is provided in the appendix. ∎
This Proposition shows the game’s properties depend on the relationship between the agent’s critical point and balance point, affected by baseline demand $X_{1},X_{2},Y_{1},Y_{2}$, CP charge parameters $\pi$, and shifting penalty parameters $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$. In the concave game conditions, all agents can’t change the CP period together, and only one period is active, determined by system baseline conditions $S_{\mathrm{b},1}$ and $S_{\mathrm{b},2}$. If all agents are capable according to Definition 3, the CP game is quasiconcave. Otherwise, when both agents can change the CP period together, but one is non-capable, the CP game is non-concave. In practical systems, due to the variant agents’ shifting capability, all three types of games are likely to appear. Empirical studies on demand response indicate that some industrial consumers with automation can shift demand efficiently, while some are inefficient in shifting [Siano, 2014]. With these properties of CP game $G$, we then analyze its NE.
III-B Nash Equilibrium
In this section, we study the NEs of the two-agent two-period CP games in all conditions as described in Proposition 4. The main theorem is as follows
Theorem 5.
NEs in two-agent two-period CP game. Given critical points $r_{\mathrm{x}},r_{\mathrm{y}}$ and balance points $b_{\mathrm{x}},b_{\mathrm{y}},b$ as defined in Definition 2, the unique pure-strategy NE $(x^{*},y^{*})$ as defined in Definition 1 of the two-agent two-period CP game $G$ satisfy one of the following:
  1. 1.
    Concave CP game.
    $x^{*}$ $=-r_{\mathrm{x}},y^{*}=-r_{\mathrm{y}},$ $\text{if }b_{\mathrm{x}}<-r_{\mathrm{x}},b_{\mathrm{y}}<-r_{\mathrm{y}}-r_{\mathrm{x}}-b_{\mathrm{x}};$ (7a)
    $x^{*}$ $=r_{\mathrm{x}},y^{*}=r_{\mathrm{y}},$ $\text{if }b_{\mathrm{x}}>r_{\mathrm{x}},b_{\mathrm{y}}>r_{\mathrm{y}}+r_{\mathrm{x}}-b_{\mathrm{x}};$ (7b)
  2. 2.
    Quasiconcave CP game.
    $x^{*}$ $=b_{\mathrm{x}},y^{*}=b_{\mathrm{y}},$ $\text{if }-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}};$ (7c)
  3. 3.
    Non-concave CP game.
    $x^{*}$ $=-r_{\mathrm{x}},y^{*}=b+r_{\mathrm{x}},$ if $\text{if }\{-r_{\mathrm{y}}-r_{\mathrm{x}}\leq b\leq r_{\mathrm{y}}-r_{\mathrm{x}}\}\cap\{b_{\mathrm{x}}<-r_{\mathrm{x}}\cup b_{\mathrm{y}}>r_{\mathrm{y}}\}$ (7d)
    $x^{*}$ $=r_{\mathrm{x}},y^{*}=b-r_{\mathrm{x}},$ if $\text{if }\{-r_{\mathrm{y}}+r_{\mathrm{x}}\leq b\leq r_{\mathrm{y}}+r_{\mathrm{x}}\}\cap\{b_{\mathrm{x}}>r_{\mathrm{x}}\cup b_{\mathrm{y}}<-r_{\mathrm{y}}\}$ (7e)
Sketch of the proof.
We comprehensively analyze the NE in all possible conditions by separating two cases according to the system CP period determined by $S_{\mathrm{b},1},S_{\mathrm{b},2}$, and further separate the cases according to agents’ own baseline demand $X_{1},X_{2},Y_{1},Y_{2}$, then study whether both agents are capable, one of them is capable, and none of them is capable. The reason is that agents’ best response is to shift demand away from the CP period when their demand is higher in the CP period, to avoid CP charges, or shift demand toward the CP period when their demand is lower in the CP period, to maintain the CP period. However, the shifting capability is determined by their parameters, i.e., the relationship between the critical point $r_{\mathrm{x}}$ and the agent balance point $b_{\mathrm{x}}$. We analyze each condition and show corresponding NE solutions. The details are provided in the appendix. ∎
This theorem shows the NEs under each concave, quasiconcave, and non-concave CP game type. Note that the condition described in this theorem aligns with the conditions in Proposition 4, and we write out individual agent’s conditions along with the NE solutions. Basically, the concave CP game includes two cases (7a) and (7b), both are aligned with (6a) and (6b) in Proposition 4. The condition of quasiconcave CP game (7c) is the same to (6c) in Proposition 4. The non-concave CP game conditions (7d) and (7e) are also aligned with (6d) in Proposition 4.
The NE also describes the relationship of gaming agents as fully cooperative (7a), (7b), fully competitive (7c), and mixed competitive and cooperative (7d), (7e). Specifically, When they are in a fully cooperative relationship, they cooperate to shift the demand away from the CP period. Once they can change the CP period together, they enter a competitive relationship where they compete with each other to change the CP period, but one agent’s decision is limited by their shifting capability. When all agents are capable, they are in a fully competitive relationship that can balance their demand regardless of opponents’ strategy. Indeed, if both agents want to be in a mixed relationship, they must be ’asymmetrical,’ where their shifting penalty parameters should be different so that they have different shifting limitations. Otherwise, they are either fully cooperative or fully competitive.
We also show in practice, under the quasiconcave and concave game conditions, agents share the market equally, i.e., their strategy only depends on their parameters, regardless of the opponent’s strategy. While in the non-concave game, the agents’ strategy depends on the opponent’s strategy, and the agent with $y=b\mp r_{\mathrm{x}}$ is more flexible and dominates the other inflexible agent with $x=\pm r_{\mathrm{x}}$ and benefit more through the competition. We denote agents’ flexibility as the demand shifting capability, either $x^{*}=\min\{r_{\mathrm{x}},b_{\mathrm{x}}\}$ or $x^{*}=\max\{-r_{\mathrm{x}},b_{\mathrm{x}}\}$. Here, $x^{*}$ depends on the game type and is influenced by the shifting penalty parameter $\alpha_{\mathrm{x}}$ and $b_{\mathrm{x}}$. Thus, flexibility is determined by both $\alpha$ and the baseline demand conditions $X_{1},X_{2},Y_{1},Y_{2}$.
IV Equilibrium stability and algorithm convergence
The CP game system is a switched dynamics system due to the indicator function in the model (1). In this section, we first analyze the system stability and then develop a solution algorithm to reach the stable (equilibrium) point.
IV-A Equilibrium stability
There are two system dynamics corresponding to periods 1 and 2, separated by the indicator function and a switching logic between these two periods. According to [Rosen, 1965], we consider a reasonable dynamic model in each period in which each agent changes his strategy following the gradient direction with respect to his strategy of his payoff function, then each agent’s payoff will increase given all other agents’ strategies. Denote the dynamic time index with $k$, and the gradient as $F_{1}(x,y)=[\nabla_{\mathrm{x}}f_{\mathrm{x}}(x,y),\nabla_{\mathrm{y}}f_{\mathrm{y}}(x,y)]^{T}$ with total differential operator $\nabla$, the dynamics of period 1 ($x,y\in\mathcal{X}_{\mathrm{cp}1}$) is
$\frac{d[x,y]^{T}}{dk}$ $=[\dot{x},\dot{y}]^{T}=F_{1}(x,y)=[-(\pi+2\alpha_{\mathrm{x}}x),-(\pi+2\alpha_{\mathrm{y}}y)]^{T}.$ (8a)
The period 2, i.e., $x,y\in\mathcal{X}_{\mathrm{cp}2}$, follow the same structure with different gradient $F_{2}(x,y)$:
$\frac{d[x,y]^{T}}{dk}$ $=[\dot{x},\dot{y}]^{T}=F_{2}(x,y)=[\pi-2\alpha_{\mathrm{x}}x,\pi-2\alpha_{\mathrm{y}}y]^{T}.$ (8b)
We then denote the CP game system with the switching logic as
$[\dot{x},\dot{y}]^{T}=\begin{cases}F_{1}(x,y)&x,y\in\mathcal{X}_{\mathrm{cp}1}\\ F_{2}(x,y)&x,y\in\mathcal{X}_{\mathrm{cp}2}\end{cases}.$ (9)
Our goal is to prove the overall CP game system with the switching logic is global uniform asymptotically stable. According to the NE described in Theorem 5, we call $(x^{*},y^{*})$ as the equilibrium point of the CP game system (9) if one of the following is satisfied,
$f_{\mathrm{x},1}(x^{*})+f_{\mathrm{y},1}(y^{*})$ $=f_{\mathrm{x},2}(x^{*})+f_{\mathrm{y},2}(y^{*}),$ (10a)
$F_{1}(x^{*},y^{*})$ $=0,S_{\mathrm{b},1}\geq S_{\mathrm{b},2},$ (10b)
$F_{2}(x^{*},y^{*})$ $=0,S_{\mathrm{b},1}<S_{\mathrm{b},2}.$ (10c)
Among them, the first condition (10a) corresponds to non-concave CP game (7d) and (7e), and quasiconcave CP game (7c), where both periods have the same CP charges and shifting penalty because the system demand is balanced in the two periods at NE and the shifting penalty is symmetric in the two periods. The second and third (10b) and (10c) conditions correspond to the concave CP game (7a) and (7b), where the NE is obtained when the gradient $F_{1}$ or $F_{2}$ reaches zero. Then, we introduce the main stability results.
Theorem 6.
Global stability of equilibrium point. The CP game system (9) is global uniform asymptotically stable in $\mathcal{X}_{\mathrm{s}}$, where
$\mathcal{X}_{\mathrm{s}}=\{x,y|\alpha_{\mathrm{x}}x^{2}+\alpha_{\mathrm{y}}y^{2}+\pi(S_{\mathrm{b},1}+x+y)>0\cap\alpha_{\mathrm{x}}x^{2}+\alpha_{\mathrm{y}}y^{2}+\pi(S_{\mathrm{b},2}-x-y)>0\},$ (11)
i.e., for every starting point $(x,y)\in\mathcal{X}_{\mathrm{s}}$, the solution $(x(k),y(k))$ to the CP game system (9) converges to an equilibrium point $(x^{*},y^{*})\in\mathcal{X}_{\mathrm{s}}$ as $k\rightarrow\infty$, where $(x^{*},y^{*})$ is the equilibrium point satisfy (10).
Sketch of the proof.
The switched dynamics system property makes the stability analysis non-trivial. We first show that the dynamical systems within each period described by (8a) and (8b) are asymptotically stable in $\mathcal{X}$. Then, we include the switching logic and derive multiple Lyapunov functions with continuous properties in the switching surface to show the CP game system (9) is global uniform asymptotically stable in $\mathcal{X}_{\mathrm{s}}$. After that, we show the stable point is obtained with $F_{1}=0$ or $F_{2}=0$ under the condition of concave CP game, and with $f_{\mathrm{x},1}+f_{\mathrm{y},1}=f_{\mathrm{x},2}+f_{\mathrm{y},2}$ under the condition of quasiconcave or non-concave CP game. From Theorem 5, we also know the equilibrium is unique, which means the stable point is the equilibrium point. The detailed proof is provided in the appendix. ∎
This Theorem shows the global uniform asymptotically stability of our CP game system in the strategy set $\mathcal{X}_{\mathrm{s}}$. Combine with Theorem 5, the CP game $G$ can converge to the unique equilibrium points for the solution trajectory that within $\mathcal{X}_{\mathrm{s}}$. It is important to know that $\mathcal{X}_{\mathrm{s}}$ is reasonable in real operations. Specifically, the key component of $\mathcal{X}_{\mathrm{s}}$ is $\pi(S_{\mathrm{b},2}-x-y)$, which means when agents shift demand, the shifting amount remains within the system’s baseline limitation. This holds naturally and ensures practical system stability, as demand shifting cannot result in negative system demand, meaning agents cannot ”borrow” demand to shift beyond their available capacity.
With the stability property, our next step is to develop an algorithm to compute the equilibrium point.
IV-B Algorithm to determine the equilibrium point
Given the finite difference approximation to the CP game system dynamics (9) with learning rate vector $\tau_{\mathrm{x},h},\tau_{\mathrm{y},h}$ for agent x and y, and gradient $F_{j}$
$[x_{h+1},y_{h+1}]^{T}=[x_{h},y_{h}]^{T}+\text{diag}(\tau_{\mathrm{x},h},\tau_{\mathrm{y},h})F_{j}(x_{h},y_{h}),j=1,2,$ (12)
where $\text{diag}(\cdot):\mathbb{R}^{2}\to\mathbb{R}^{2\cdot 2}$, $j$ is the switching signal taking the value 1 for $S_{1}(x,y)\geq S_{2}(x,y)$ and 2 for $S_{1}(x,y)<S_{2}(x,y)$. This then forms a gradient-based algorithm following the updating rule (12) to gradually reach the equilibrium point. Among them, the gradient is determined by the period on which the current solution lies, and the learning rate for agents x and y $\tau_{\mathrm{x},h},\tau_{\mathrm{y},h}$ are determined by its payoff function and gradient. Choosing a suitable learning rate is key to showing convergence performance, we thus provide the following Theorem.
Theorem 7.
Determination of the equilibrium point. Given the finite difference approximation as described in (12), a finite learning rate $\tau_{\mathrm{x},h}$ for agent x can be selected such that when $X_{1}+x\geq X_{2}-x$
$-f_{\mathrm{x},1}(x_{\overline{h}})<-f_{\mathrm{x},1}(x_{\underline{h}}),-f_{\mathrm{x},2}(x_{\overline{h}})>-f_{\mathrm{x},2}(x_{\underline{h}}),$ (13a)
when $X_{1}+x<X_{2}-x$,
$-f_{\mathrm{x},1}(x_{\overline{h}})>-f_{\mathrm{x},1}(x_{\underline{h}}),-f_{\mathrm{x},2}(x_{\overline{h}})<-f_{\mathrm{x},2}(x_{\underline{h}}),$ (13b)
where $F_{j}(x_{h},y_{h})\neq 0,j=1,2$; $\underline{h}$ and $\overline{h}$ is a switched pair that satisfies $\underline{h}<h<\overline{h}$ and $\underline{h}=\overline{h}=j$, $h\neq j$. The same also holds for the learning rate $\tau_{\mathrm{y},h}$ of agent y.
Sketch of the proof.
We use the backtracking line search method [Boyd and Vandenberghe, 2004] to calculate the learning rate, which is affected by the CP periods of the current and future steps. For the concave game, only one period could be the CP period during the entire solution process, according to the backtracking line search, the gradient $F_{1}$ or $F_{2}$ reduce gradually, and combined with Theorem 6, the gradient will reduce to zero and reach the equilibrium point satisfy (10b) or (10c).
When switching happens, agents’ individual peak periods are different, and we analyze the objective function $-f_{\mathrm{x},1},-f_{\mathrm{x},2}$ change following the backtracking line search criteria by judging where the CP periods is in the current and next steps. Given a trajectory starts from period 1 at $h$, switch to period 2 at $h+1$, and back to period 1 at $h+2$, suppose agent x’s individual peak period is 1, we then show agent x updates its decision $x$ when switching from period 1 to 2, and agent y update its decision $y$ when switching from period 2 to 1, i.e., the $\tau_{\mathrm{y},h}=0,\tau_{\mathrm{x},h+1}=0$, correspondingly, agent x’s objective $-f_{\mathrm{x},1}$ decrease and agent y’s objective $-f_{\mathrm{y},2}$ decrease. The reason for the change is that their individual peak period aligns with the system’s CP period. Note that this learning rate is determined by their payoff functions and gradients, allowing the gradient for agent x to increase its decision variable while the gradient for agent y keeps the decision variable. By the same analysis, we know agent x’s objective $-f_{\mathrm{x},2}$ increase and agent y’s objective $-f_{\mathrm{y},1}$ increase in the trajectory starting from period 2, switched to period 1, and back to period 2. We then show these results still hold when the trajectory stays more steps in the system between two transitions. This proves the reduction of the distance between objective functions in the two periods. Combined with Theorem 6, the entire system finally will satisfy (10a) and reach the equilibrium point. The details are provided in the appendix. ∎
This Theorem shows that the finite learning rate can be chosen in each step using the backtracking line search. Following this updating rule (12), combined with Theorem 6, a gradient-based algorithm can reach the equilibrium point. Specifically, the difference in each agent’s payoff function in the two periods reduces gradually, and depending on whether the game is concave or not, the difference can be reduced to zero or until one of the periods’ gradients reaches zero.
As Theorem 6 shows, the system is asymptotically stable in $\mathcal{X}_{\mathrm{s}}$. Combined with Theorem 7, this confirms that the algorithm computes the NE as described in Theorem 5. When applying this game framework to a practical CP shaving system, the convergence process naturally reflects real agent interactions in the gaming process. Specifically, the system dynamics alternate between periods 1 and 2, while agents iteratively update their decisions based on the best response strategy. That is, given the opponents’ (aggregated) strategy, each agent determines an optimal response based on its own payoff. This iterative process follows the system dynamics and ultimately converges to the NE described in Theorem 5. Note that we don’t use a higher-order gradient descent method, such as Newton’s method because the fast updating with higher-order gradient information lets the solution in each period converge to each period’s stable point too fast to realize the converge on the overall switched system.
V Impact of Customers Strategic Behavior
In this section, we analyze gaming agents’ strategic behavior in the two-agent two-period setting. First, we state a benchmark centralized CP shaving model for comparison, then analyze agents’ strategic behavior from both an economic perspective with the PoA, and a technical perspective with the peak shaving effectiveness.
V-A Centralized CP shaving
We first state the centralized CP shaving model, which assumes a central operator has direct control over both agents in the formulated two-period condition to reduce the total cost of both agents. This centralized model thus maximizes the total objective function of both agents and is denoted as
$(x^{*},y^{*})\in\arg\max_{x,y}f_{\mathrm{x}}(x,y)+f_{\mathrm{y}}(x,y),$ (14)
We now present the following proposition, demonstrating that the centralized CP shaving model is equivalent to the centralized peak shaving model. This model minimizes the total peak demand of both agents while also incorporating peak demand charges and shifting costs.
Proposition 8.
Centralized peak shaving model. The centralized CP shaving model (14) is equivalent to the trivial convex peak shaving model as follows:
$(x^{*},y^{*})\in\arg\min_{x,y}\ \pi\max\{S_{1}(x,y),S_{2}(x,y)\}+\alpha_{\mathrm{x}}x^{2}+\alpha_{\mathrm{y}}y^{2}$ (15)
Proof.
Take $f_{\mathrm{x}},f_{\mathrm{y}}$ as defined in (1) into (14), we have
$f_{\mathrm{x}}(x,y)$ $+f_{\mathrm{y}}(x,y)=\pi(X_{1}+x)I(S_{1}(x,y)-S_{2}(x,y))+\pi(X_{2}-x)I(S_{2}(x,y)-S_{1}(x,y))$
$+\pi(Y_{1}+y)I(S_{1}(x,y)-S_{2}(x,y))+\pi(Y_{2}-y)I(S_{2}(x,y)-S_{1}(x,y))+\alpha_{\mathrm{y}}y^{2}+\alpha_{\mathrm{x}}x^{2}$
$=\pi S_{1}(x,y)I(S_{1}(x,y)-S_{2}(x,y))+\pi S_{2}(x,y)I(S_{2}(x,y)-S_{1}(x,y))+\alpha_{\mathrm{y}}y^{2}+\alpha_{\mathrm{x}}x^{2}$
$=\pi\max\{S_{1}(x,y),S_{2}(x,y)\}+\alpha_{\mathrm{y}}y^{2}+\alpha_{\mathrm{x}}x^{2}.$ (16)
Note that the peak shaving model in (15) is convex because maximize two convex (linear) functions $S_{1}(x,y),S_{2}(x,y)$ is convex; thus, we prove the Proposition. ∎
This model assumes direct control over each agent’s demand, achieving peak shaving at minimal cost. However, while its performance is strong, the assumption of direct load control is unrealistic in practice, as agents (customers) typically have independent strategies and privacy concerns. Therefore, we consider (15) as a benchmark to compare with our CP game model and analyze peak shaving effectiveness and the PoA in the following sections.
V-B Peak shaving ratio analysis
We show in the following theorem that the CP game can achieve the same effectiveness in reducing the system peak demand at its equilibrium.
Theorem 9.
Peak shaving performance. The peak shaving effectiveness of the CP game model $G$ at equilibrium is always 1, i.e.,
$\frac{\max\{S_{1}(x^{*},y^{*}),S_{2}(x^{*},y^{*})\}}{\max\{S_{1}(x_{\mathrm{cen}}^{*},y_{\mathrm{cen}}^{*}),S_{2}(x_{\mathrm{cen}}^{*},y_{\mathrm{cen}}^{*})\}}=1,$ (17)
for all $\pi,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}},X,Y>0$; where $x^{*},y^{*}$ is the game equilibrium results and $x_{\mathrm{cen}}^{*},y_{\mathrm{cen}}^{*}$ is the centralized peak shaving results.
Sketch of the proof.
We prove this Theorem by first analytically writing the solution for the centralized model shown in Proposition 8, and from Theorem 5, we have the game model solution. Then take the solutions of the centralized model and the game model under the same condition into the peak shaving effectiveness definition in this Theorem. The detailed proof is provided in the appendix. ∎
This Theorem shows the game model always reaches the same peak shaving performance when compared with the centralized model. The reason is that agents in both the game model and centralized model shift demand as much as possible to avoid CP charge without burdening more by their shifting penalty. Specifically, under concave game conditions, both agents’ shifting capability is constrained by their critical points, preventing them from fully balancing their demand—similar to the centralized model. Under quasiconcave and non-concave game conditions, system demand is balanced across the two periods, also aligning with the centralized model. This suggests that the game model is practically applicable, as utilities and operators can implement the game framework for CP shaving without compromising peak shaving performance. In other words, operators can still achieve their peak shaving targets while using the game model. However, although the overall peak shaving performance is the same, agents’ individual demand shifting is different due to the information barrier, which reflects as cost of reaching the peak shaving performance. We then analyze the cost by showing the PoA in the following section.
V-C PoA analysis
In this section, we study the PoA affected by gaming agents’ strategic behavior in all three game structures, and we provide the following main results.
Theorem 10.
PoA with agent equity. Given the PoA defined as
$P=\frac{f_{\mathrm{x}}(x^{*},y^{*})+f_{\mathrm{y}}(x^{*},y^{*})}{f_{\mathrm{x}}(x_{\mathrm{cen}}^{*},y_{\mathrm{cen}}^{*})+f_{\mathrm{y}}(x_{\mathrm{cen}}^{*},y_{\mathrm{cen}}^{*})},$ (18a)
under the quasiconcave and non-concave game condition as described in Proposition 4, the PoA increases with the disparities among agents, as measured by the marginal shifting cost $\alpha_{\mathrm{x}}x^{*},\alpha_{\mathrm{y}}y^{*}$, i.e.,
$\frac{\partial P}{\partial[(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}]}>0$ (18b)
where $x^{*},y^{*}$ is the game equilibrium result and $x_{\mathrm{cen}}^{*},y_{\mathrm{cen}}^{*}$ is the centralized peak shaving results.
Sketch of the proof.
From Theorem 5, we know the NE of the game model under quasiconcave and non-concave game conditions. Combined with the centralized model solution obtained from Theorem 9, we show the difference of nominator and denominator of the PoA as defined in (18a) can be expressed as a Euclidean distance between agents’ marginal shifting cost, i.e., $(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}$. The detail is provided in the appendix. ∎
Under non-concave and quasiconcave game conditions, although the overall CP charge is always the same as the centralized model due to the same peak shaving effectiveness, the shifting cost $\alpha_{\mathrm{x}}x^{*2}+\alpha_{\mathrm{y}}y^{*2}$ increases due to the information barrier. We show in this theorem that the PoA increases with agents’ disparities, quantified by their marginal shifting cost. This insight provides a pathway for designing mechanisms that balance agents’ marginal shifting costs to ensure both system effectiveness and fairness. For instance, agents with both a large shifting penalty parameter and a high shifting amount bear a higher marginal shifting cost. Here, the penalty parameter reflects greater comfort loss from demand shifting, while the high shifting amount arises from significant demand differences requiring adjustment due to CP charges. In such cases, operators or utilities could regulate these agents to shift less demand or provide incentives to reduce their penalty parameters. As some policies already suggest subsidizing disadvantaged customers, this analysis offers guidance on how to allocate subsidies effectively among customers [Pacific Gas & Electric Company, 2025, Consolidated Edison, 2025b]. By balancing marginal shifting costs, this approach simultaneously improves system equity and reduces the PoA.
Theorem 11.
PoA with CP game type. For given $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$, the PoA defined as Theorem 10 is highest at equilibrium for quasiconcave games, followed by non-concave games, and lowest for concave games, where it is always equal to 1, i.e.,
$P(\text{Quasiconcave game})\geq P(\text{Non-concave game})\geq P(\text{Concave game})=1.$
Sketch of the proof.
We first show $P=1$ is always true for the concave game due to the same solution structure between the centralized model and concave game model based on the results from Theorems 5 and 9. Then from Theorem 10, we know the PoA expression as Euclidean distance between agents’ marginal shifting cost under quasiconcave and non-concave game conditions, indicating $P\geq 1$. Thus, quasiconcave games and non-concave games always cause higher (or equal) PoA than concave games. We then show by fixing $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$, the agent’s baseline demand $X_{1},X_{2},Y_{1},Y_{2}$ can vary to cause different game structures, so as to different solution structure $x^{*},y^{*}$ from Theorems 5. Combined with the fact that quasiconcave and non-concave games always balance system demand, i.e., $x^{*}+y^{*}=b$ from Theorem 5, we prove the quasiconcave game solution shows a higher difference between $x^{*},y^{*}$, indicating higher PoA than a non-concave game. The detailed proof is provided in the appendix. ∎
This theorem demonstrates the impact of the CP game type on the PoA, where the game type connects to the agents’ flexibility. We first fix the system conditions $\pi,S$ while allowing the agents’ conditions to vary. From Theorem 5, we know that agents’ flexibility is determined by their shifting penalty parameters $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$ and baseline demand $X_{1},X_{2},Y_{1},Y_{2}$. To control the influence of both parameters, we also fix $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$, making the PoA dependent only on $x^{*},y^{*}$, which reflects the difference in the game structure. As the balance point (demand difference) $b_{\mathrm{x}}=(X_{2}-X_{1})/2$ increases, agent x’s flexibility decreases, and the game type transitions from a quasiconcave game to a non-concave game, and finally to a concave game. This follows from the conditions $\{-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}\}\cap\{-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}}\}$ to $\{-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}\}\cap\{b\geq r_{\mathrm{x}}+r_{\mathrm{y}}\cup b\leq-r_{\mathrm{x}}-r_{\mathrm{y}}\}$ to $\{r_{\mathrm{x}}+r_{\mathrm{y}}<b\}\cup\{-r_{\mathrm{x}}-r_{\mathrm{y}}>b\}$ as stated in Proposition 4. Thus, the PoA increases with agents’ flexibility and is also reflected in the game type change. This highlights the role of inflexible agents in overall system performance, despite utilities and operators generally favoring flexible agents for their ability to enhance grid stability, integrate renewable energy, and respond to pricing signals [Siano, 2014]. It is also evident that fixing $X_{1},X_{2},Y_{1},Y_{2}$ while allowing $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$ to vary yields the same results. Note that this theorem also shows that under the concave game condition, the system’s performance is always equivalent to that of the centralized model, indicating no harm to utility companies and agents when applying the concave CP game.
VI Extension to Multi-agent Games
In this section, we consider a generalization of our model from the two-agent two-period CP game to the multi-agent two-period CP game, and we analyze the NE, stability/convergence, as well as the consumers’ strategic behavior.
VI-A Multi-agent CP game
We denote $X_{i,1},X_{i,2}$ as agent $i$’s baseline demand in period 1 and 2, $\alpha_{i}$ as agent $i$’s penalty parameter, and $x_{i}$ as agent $i$’s demand shifting strategy. We then extend the CP game $G$ from two-agent to multi-agent with $G^{\prime}=(N,\mathcal{X},U)$, such that,
  • $N=\{1,2,...,|N|\}$ with index $i$ is the agent set, where ||$\cdot$|| ⋅ | for a set means the number of elements inside a set (otherwise, it is absolute value); specifically, we use $-i$ to indicate all agents but $i$.
  • $\mathcal{X}=\times_{i\in N}\mathcal{X}_{i}$ is the strategy set formed by the product topology of each agent’s individual strategy set $\mathcal{X}_{i}$.
  • $U=\{f_{i}|i\in N\}$ is the payoff function set, where $f_{i}:\mathcal{X}\rightarrow\mathbb{R}$.
Then agent $i$’s payoff function as described in (1) can be generalized as follows:
$\max_{x_{i}}$ $f_{i}(x_{i},x_{-i})=-\pi(X_{i,1}+x_{i})I(S_{1}(x_{i},x_{-i})-S_{2}(x_{i},x_{-i}))$
$-\pi$ $(X_{i,2}-x_{i})I(S_{2}(x_{i},x_{-i})-S_{1}(x_{i},x_{-i}))-\alpha_{i}x_{i}^{2},$ (19a)
$S_{1}(x_{i},x_{-i})$ $=\sum_{i\in N}(X_{i,1}+x_{i})=S_{\mathrm{b},1}+\sum_{i\in N}x_{i},$ (19b)
$S_{2}(x_{i},x_{-i})$ $=\sum_{i\in N}(X_{i,2}-x_{i})=S_{\mathrm{b},2}-\sum_{i\in N}x_{i},$ (19c)
Note that the system demands $S_{1},S_{2}$ change from the summation of two agents’ to $N$ agents’ demand at periods 1 and 2.
The multi-agent CP game $G^{\prime}$ still follow the property as described in Proposition 4, i.e., the CP game $G^{\prime}$ is
  1. 1.
    Concave if
    $b<-\sum_{i\in N}r_{i},\text{when }S_{\mathrm{b},1}\geq S_{\mathrm{b},2},b>\sum_{i\in N}r_{i},\text{when }S_{\mathrm{b},1}<S_{\mathrm{b},2},$ (20a)
  2. 2.
    Quasiconcave if
    $-r_{i}\leq b_{i}\leq r_{i},\forall i\in N;$ (20b)
  3. 3.
    Non-concave if
    $\{-\sum_{i\in N}r_{i}\leq b\leq\sum_{i\in N}r_{i}\}\cap\{\exists i\in N,b_{i}<-r_{i}\cup b_{i}>r_{i}\}.$ (20c)
The concave and quasiconcave conditions are intuitive and can be obtained following the same process of Proposition 4. Then the non-concave condition is determined by the complementary set of concave and quasiconcave conditions.
In terms of the NE solutions, we can’t analyze each agent’s capability case by case like the two-agent setting. Thus, we separately analyze NE in each game type. We first show concave game has a unique NE according to [Rosen, 1965], where the NE definition slightly changes from Definition 1 to include all agents, i.e.,
$f_{i}(x_{i}^{*},x_{-i}^{*})\geq f_{i}(x_{i},x_{-i}^{*}),i\in N,x_{i}\in\mathcal{X}_{i},x_{-i}\in\mathcal{X}_{-i}.$ (21)
The NE is then obtained when all agents reach their critical point as defined in Definition 2 corresponding to the active periods, and we have the NE as follows,
$x_{i}^{*}$ $=-r_{i},i\in N,\text{if }b<-\sum_{i\in N}r_{i},$ (22a)
$x_{i}^{*}$ $=r_{i},i\in N,\text{if }b>\sum_{i\in N}r_{i}.$ (22b)
Noted that when all $x_{i}^{*}$ take $-r_{i}$, the CP period is always period 1 and $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$, and agents shift demand from period 1 to period 2, until reaching their critical point.
NE of the quasiconcave game is not intuitive and we introduce the following Proposition to analyze it.
Proposition 12.
Existence and uniqueness of NE in multi-agent quasiconcave CP game. The quasiconcave CP game $G^{\prime}$ as described in (19) and satisfy (20b) has a unique pure-strategy NE $(x_{i}^{*},x_{-i}^{*})$ as defined in (21), where
$x_{i}^{*}=b_{i},i\in N,S_{1}(x_{i}^{*},x_{-i}^{*})=S_{\mathrm{b},1}+\sum_{i\in N}x_{i}^{*}=S_{\mathrm{b},2}-\sum_{i\in N}x_{i}^{*}=S_{2}(x_{i}^{*},x_{-i}^{*}).$ (23)
Sketch of the proof.
To prove this Proposition, we derive two Lemmas showing the existence (Lemma 17) and uniqueness (Lemma 18). The basic principle of existence Lemma is based on [Reny, 1999], i.e., the sum of the agent’s payoff functions is upper semi-continuous in $x_{i}\in\mathcal{X}$ and game $G^{\prime}$ is payoff security, where we define in the proof. The proof of uniqueness Lemma is based on [Rosen, 1965]. We first prove the uniqueness of NE in periods 1 and 2. Then, we start with a two-agent setting to analyze all agents’ best responses and use a min-max formulation to show the NE of each agent. After that, we apply these two-agent results to multi-agent settings by sequentially partitioning agents into two groups. The detailed proof is provided in the appendix. ∎
This Proposition shows the quasiconcave CP game $G^{\prime}$ has a unique NE and always obtains at their balance points $b_{i}$. Note that under both concave and quasiconcave conditions, the NE results remain consistent with the two-agent setting, as the extension from two agents to multiple agents follows a straightforward generalization. This suggests an approach where the NE results of the two-agent non-concave CP game $G$ serve as a foundation for analyzing the $N$-agent non-concave CP game $G^{\prime}$, whose NE is evidently non-trivial. Basically, Theorem 5 conveys an important message that the system demand will always be balanced in the two periods, i.e., $x^{*}+y^{*}=b$, and one agent (less flexible according to shifting penalty parameters and baseline demand) first reach the critical point while the other agent balance the system demand in the two periods.
Here we denote the agents whose baseline peak demand is in the system baseline CP period as CP-period agent, and the other as non-CP-period agent. We also call the set that includes all CP-period agents as CP-period agent set and includes all non-CP-period agents as non-CP-period agent set, denote as $N_{\mathrm{cp}},N_{\mathrm{ncp}}$, respectively, such that $N_{\mathrm{cp}}\cup N_{\mathrm{ncp}}=N,N_{\mathrm{cp}}\cap N_{\mathrm{ncp}}=\emptyset$. We then provide the following Proposition to analyze the equilibrium condition of the non-concave multi-agent CP game $G^{\prime}$.
Proposition 13.
NE in non-concave multi-agent CP game. Given non-concave multi-agent CP game $G^{\prime}$ and $S_{\mathrm{b},1}<S_{\mathrm{b},2}$, under the condition of
$\{\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}<\sum_{i\in N_{\mathrm{cp}}}b_{i}\}\cap\{-\sum_{i\in N}r_{i}\leq b\leq\sum_{i\in N}r_{i}\};,$ (24a)
the hybrid NE solution, i.e., for each CP-period agent and for the whole non-CP-period agent set, is given by
$x_{i}^{*}=\min\{r_{i},b_{i}\},i\in N_{\mathrm{cp}},\sum_{i\in N_{\mathrm{ncp}}}x_{i}^{*}=b-\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\};$ (24b)
otherwise, the hybrid NE solution, i.e., for each non-CP-period agent and the overall CP-period agent set, is given by
$x_{i}^{*}=\max\{-r_{i},b_{i}\},i\in N_{\mathrm{ncp}},\sum_{i\in N_{\mathrm{cp}}}x_{i}^{*}=b-\sum_{i\in N_{\mathrm{ncp}}}\max\{-r_{i},b_{i}\},$ (24c)
corresponding to the condition
$\{\sum_{i\in N_{\mathrm{ncp}}}\max\{-r_{i},b_{i}\}>\sum_{i\in N_{\mathrm{ncp}}}b_{i}\}\cap\{-\sum_{i\in N}r_{i}\leq b\leq\sum_{i\in N}r_{i}\};$ (24d)
Sketch of the proof.
The basic idea is to assume two virtual agents that have the same baseline demand and shifting penalty conditions with the CP-period agent set and non-CP-period agent set. However, their strategy structures are different, where the agent set’s strategy is determined by the aggregation of each individual agent’s strategy. From Theorem 5, we derive the best strategies of two virtual agents, and from the best response analysis, we know the agent sets and virtual agents have the same response rationale due to the same baseline conditions. Basically, they will both shift demand away from the CP period or to the CP period to avoid the cost, and the cost reduction positively relates to their shifting amount. We then link the two virtual agents’ strategies to the agent sets’ strategy and show the equilibrium point for the CP-period agent set and the non-CP-period agent set under different baseline conditions. The detailed proof is provided in the appendix. ∎
In the Proposition, we present the condition for $S_{\mathrm{b},1}<S_{\mathrm{b},2}$; when $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$, the only difference is to replace ±$\pm r_{i}$ with $\mp r_{i}$. This Proposition demonstrates that we can analytically determine the equilibrium solution for one set of agents — either the CP-period agent set $N_{\mathrm{cp}}$ or non-CP-period agent set $N_{\mathrm{ncp}}$. For the remaining set, we can determine only the aggregated performance of the set. Thus, we call it a hybrid NE, where one subset of agents follows a strict NE, while another subset allows internal flexibility while maintaining the set-level balance. Essentially, more flexible agents, those with lower shifting penalty parameters and higher baseline demand differences, will shift more to achieve the overall shifting performance of their set. One important insight from this proposition is that system demand remains balanced under non-concave game conditions, aligning with the two-agent setting analysis. This result helps in evaluating system performance in terms of PoA and peak shaving effectiveness.
We then analyze the stability of the equilibrium point in the multi-agent CP game $G^{\prime}$ based on the two-agent game results presented in Theorem 6. Extending the gradient vector to include all agents’ shifting decision variables $x_{i},i\in N$, we still denote the system dynamics as (9). Following the same proof structure, we can derive the multiple Lyapunov functions for periods 1 and 2, demonstrating that the global uniform asymptotically stable property also holds for the multi-agent CP game $G^{\prime}$. Under concave and quasiconcave conditions, the unique NE described in (22) and Proposition 12 serve as the stable equilibrium point in the system dynamics. In the non-concave condition, the system dynamics in periods 1 and 2 converge to the switching surface where the payoffs of both periods are equal, ensuring system demand remains balanced, as characterized by the hybrid NE in Proposition 13. The strategy set $\mathcal{X}_{\mathrm{s}}$ change to include all agents’ decisions as follows
$\mathcal{X}_{\mathrm{s}}=\{x_{i}|\sum_{i\in N}\alpha_{i}x_{i}^{2}+\pi(S_{\mathrm{b},1}+\sum_{i\in N}x_{i})>0\cup\sum_{i\in N}\alpha_{i}x_{i}^{2}+\pi(S_{\mathrm{b},2}-\sum_{i\in N}x_{i})>0\}.$ (25)
With this stability property, we then show that a finite learning rate can still be chosen such that a gradient-based algorithm using (12) as an updating rule finds the equilibrium point for the multi-agent CP game $G^{\prime}$. The only difference in the multi-agent setting is that the criteria shift from the convergence of each agent’s payoff across two periods to the convergence of the aggregated payoff of all CP-period agents and non-CP-period agents across the two periods. This also demonstrates that, under non-concave conditions, agents can internally adjust their demand-shifting strategies while ensuring that system demand remains balanced across the two periods.
Remark 14.
Determination of equilibrium point in multi-agent CP game. A finite learning rate vector $\tau_{h}\in\mathbb{R}^{N}$ of a finite difference approximation to a system dynamics in the multi-agent CP game setting can be chosen such that when $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$,
$\sum_{i\in N_{\mathrm{cp}}}-f_{i,1}(x_{i,\overline{h}})<\sum_{i\in N_{\mathrm{cp}}}-f_{i,1}(x_{i,\underline{h}}),\sum_{i\in N_{\mathrm{cp}}}-f_{i,2}(x_{i,\overline{h}})>\sum_{i\in N_{\mathrm{cp}}}-f_{i,2}(x_{i,\underline{h}}).$ (26a)
$\sum_{i\in N_{\mathrm{ncp}}}-f_{i,2}(x_{i,\overline{h}})<\sum_{i\in N_{\mathrm{ncp}}}-f_{i,2}(x_{i,\underline{h}}),\sum_{i\in N_{\mathrm{ncp}}}-f_{i,1}(x_{i,\overline{h}})>\sum_{i\in N_{\mathrm{ncp}}}-f_{i,1}(x_{i,\underline{h}}).$ (26b)
When $S_{\mathrm{b},1}<S_{\mathrm{b},2}$, Greater than sign and less than sign reverses.
To conclude, the game framework is also workable for the multi-agent two-period setting as the (hybrid) NE still exists and is global uniform asymptotically stable within a strategy set as defined in (25). Also, a gradient-based algorithm with the updating rule that approximates the system dynamics can still compute the equilibrium point. We then analyze the impact of consumers’ strategic behavior in the multi-agent setting.
VI-B Consumers strategic behavior in Multi-agent CP game
We show the peak shaving effectiveness and PoA under the multi-agent setting in this section, where the centralized model could be denoted as $(x_{i,\mathrm{cen}}^{*},x_{-i,\mathrm{cen}}^{*})\in\arg\max_{x_{i},x_{-i}}\sum_{i\in N}-f_{i}(x_{i},x_{-i})$. Combined with Proposition 8, under the concave game condition, the centralized solution can be easily obtained by the first-order optimality conditions.
The peak shaving effectiveness, as defined in Theorem 9 with the extension from agent y to $-i$ denoting all agents but $i$, is still equal to 1 at the equilibrium of the game model. The reason is that the system demand will be balanced under all game conditions, the same as the centralized model. The PoA in the multi-agent setting is $P_{N}=\sum_{i\in N}f_{i}(x_{i}^{*},x_{-i}^{*})/\sum_{i\in N}f_{i}(x_{\mathrm{cen},i}^{*},x_{\mathrm{cen},-i}^{*})$. As the centralized model is convex from Proposition 8, and $x_{\mathrm{cen},i}^{*},x_{\mathrm{cen},-i}^{*}$ is the unique minimizer for the problem, we always have $P\geq 1$. Specifically, the concave game model remains equivalent to the corresponding centralized model and always has $P=1$.
With the agent number increase, the PoA under quasiconcave and non-concave games will be affected due to the game type changes, and we have the following remark.
Remark 15.
Game type with agent numbers. As $N$ increases, the game structure will more likely be a non-concave game.
It is intuitive that the non-concave game type is more likely to appear as the agent number increases because the probability that all agents are capable or non-capable reduces exponentially. As we analyzed at Theorem 11, the agent’s flexibility changes game types and thus causes different PoA. This means the PoA of a small system is more sensitive to the agents’ flexibility, and a large system can diminishing flexible agents influence on the system efficiency. This suggests a strategy for operators and utilities when implementing a game-based CP shaving program: selecting more flexible customers to form a large CP shaving system while grouping a few inflexible agents into a smaller CP shaving system. These two systems would operate separately, with individual CP periods determined by the customers within each system. This approach enhances overall system efficiency.
VII Numerical Example
In this section, we use numerical simulations to show the CP game solution, and we show the numerical test aligns with our theoretical analysis. We set the CP charge price to $\pi=1$ and set three two-agent two-period cases and one multi-agent two-period case to show the CP game performance. In the multi-agent setting, we also test the influence of agents’ numbers on PoA.
VII-A Two-agent two-period CP game
We set the baseline demand as $X_{1}=6,X_{2}=3,Y_{1}=3,Y_{2}=10$, then change the shifting penalty parameters to change the agent capability.
(1) Set $\alpha_{\mathrm{x}}=0.2,\alpha_{\mathrm{y}}=0.1$, then all agents’ are capable according to Definition 3 and the game is quasiconcave. We show the solution path in Fig. 1, where the system alternates between periods 1 and 2. With each transition back to the previous period, the cost either decreases or increases in a way that reduces the cost difference for each agent between the two periods. Finally, the cost distance reduces to zero for each agent as they balance their demand over the two periods, and the system converges to the equilibrium point, corresponding to (7c). Compared with the solution from the centralized model, we observe that there are huge shifting changes and an increase in the system’s overall cost due to anarchy, reflected as the PoA is 1.125, indicating anarchy increases the system cost by 12.5% compared to the centralized method. The peak shaving ratio is the same as they all balance system demand.
Figure 1: Convergence of two-agent quasiconcave CP game.
(2) Set $\alpha_{\mathrm{x}}=0.5,\alpha_{\mathrm{y}}=0.1$, then agent x is non-capable, and agent y is capable according to Definition 3, and the game is non-concave. We show the solution trajectory in Fig. 2, which is similar to the quasiconcave condition. However, each agent’s cost difference over the two periods does not reach zero upon convergence, as neither agent fully self-balances its demand. Instead, the more flexible agent (agent y) achieves greater cost reduction by lowering its demand during the CP period. The results align with (7d). Compared with the quasiconcave game, we observe that the demand shifting from the game model is closer to the centralized solution, and the system cost increase due to anarchy is also reduced with a PoA of 1.0941.
Figure 2: Convergence of two-agent non-concave CP game.
(3) Set $\alpha_{\mathrm{x}}=0.5,\alpha_{\mathrm{y}}=0.6$, then both agents can’t change the CP period together, and the game is concave, where only period 2 is active during the entire solution trajectory. Thus, as we show in Fig. 3, the trajectory doesn’t jump and gradually reduces to the critical point, which corresponds to (7b). Obviously, the demand shifting converges to the same point with the centralized model, and the PoA is 1, indicating these two models are exactly equivalent.
Figure 3: Convergence of two-agent concave CP game.
VII-B Multiple agents and two-period CP game
We create a six-agent system with baseline demand and shifting penalty parameters, as Table I shows. We first notice the game is non-concave because all agents can change CP period together, but not all agents are capable, i.e., agents 3 and 4 are non-capable. Also, we know agents 1, 3, and 6 are non-CP-period agents, and agents 2, 4, and 5 are CP-period agents.
Fig. 4 shows the convergence and cost trajectory, and Fig. 5 shows the cost trajectory of non-CP period agent set and CP period agent set. Noted that the results align with our analysis in Proposition 13. Specifically, non-CP period agents shift demand either to the balance point or until they reach their critical limit. For example, agents 1 and 6 balance their demand across the two periods, while agent 3 reaches its shifting limit. After approximately 120 iterations, system demand balances over two periods (Figure 5), and CP-period agents begin internally adjusting their demand. Notably, agents 4 and 5, which have higher shifting penalty parameters than agent 2, gradually reduce their demand shifting, while agent 2 increases its shifting. From the cost trajectory (Figure 4), agent 4, with the highest shifting penalty, reduces its demand shifting more significantly than agent 5. Compared to the centralized model, the CP game equilibrium deviates significantly, with all agents shifting more demand due to the information barrier. This increases the cost and results in a PoA of 1.1317 while maintaining the same peak shaving performance.
TABLE I: Agents’ parameters and solution of game and centralized model
Agent 1 2 3 4 5 6 Total
Baseline demand 1 7 3 10 1 2 5 28
Baseline demand 2 3 13 4 4 6 3 33
Penalty parameter 0.2 0.1 0.4 0.5 0.2 0.1 \
Centralized shifting 0.36 0.72 0.18 0.15 0.36 0.73 2.5
Game shifting -2 3.85 -1.25 0.93 1.97 -1 2.5
PoA 1.1317
Figure 4: Convergence of multi-agent concave CP game.
Figure 5: Convergence of CP period agent and non-CP period agent.
VII-C Agent number impacts on PoA
In this section, we randomly generate agent samples for the systems with different agent numbers while satisfying the non-concave and quasiconcave game conditions as described in (20b) and (20c), i.e., $\sum_{i\in N}-r_{i}<b<\sum_{i\in N}r_{i}$. We set the agent’s $i,i\in N$ baseline demand as $X_{i,1},X_{i,2}\in(0,15)$ and penalty parameters as $\alpha_{i}\in(0,0.5)$. We loop the agent number from 2 to 50, and each agent number generates 1000 samples to calculate the PoA. We present the results in Fig. 6. As our theoretical analysis indicates, PoA is more fluctuates when the agent number is small; also, the game type is more varied with quasiconcave, concave, and non-concave, affected by agents’ flexibility, indicating the small system is more sensitive to the agent’s flexibility. When the agent number increases, all game conditions become non-concave, the mean PoA converges, and the variance decreases. This means the large system is more stable and can eliminate the agent’s flexibility influence on PoA.
Figure 6: Agent number impacts on PoA. Each box contains 1000 samples, the circle denotes the outlier, the box upper and lower line denotes the quartiles, and the short blue line inside the box is the median value.
VIII Discussion and Conclusion
Although CP shaving has become a common practice, the interactive influence of agents on CP time remains largely unexplored. This paper takes a first step toward understanding the application of the game framework in the context of CP shaving. We propose a theoretical game framework and show its applicability for the CP shaving problem by analyzing the game structure and analytically deriving the NE for each game structure under a two-agent two-period setting. We prove the equilibrium points are globally uniformly asymptotically stable, provided that all agents’ demand shifting remains within the baseline limit. We also show that the gradient-based algorithm, with an updating rule that approximates the asymptotically stable process via finite differences, can efficiently compute the equilibrium points. Our analysis reveals that agents’ strategic interactions in the game framework achieve the same peak-shaving effectiveness as an equivalent centralized peak-shaving model. However, under quasiconcave and non-concave conditions, the CP game results in a higher PoA. Our findings also persist for a general setting with more than two agents. Our insights on peak shaving effectiveness have the potential to inform utilities and operators that are concerned with applying game models.
By analyzing the PoA in quasiconcave and non-concave games within the two-agent setting, we show that PoA increases with the disparities between agents, as measured by their marginal shifting costs. This insight informs the design of CP shaving mechanisms that simultaneously achieve effectiveness and fairness by aligning marginal shifting costs - either by incentivizing non-flexible agents with higher comfort loss or regulating them to shift less demand. We show that PoA increases as the game transitions from concave to non-concave and then to quasiconcave, corresponding to greater agent flexibility. This suggests that higher flexibility amplifies system inefficiency. Combined with our multi-agent analysis, which shows that PoA is sensitive to agent flexibility in small systems but stabilizes in larger ones, we suggest that forming large systems with flexible agents while grouping inflexible agents into smaller systems can effectively reduce inefficiency.
Our work has a few limitations. First of all, we formulate the demand-shifting penalty as a quadratic function. Future research is needed to examine more generic functions and even obtain empirical models from data-driven methods. Another limitation is that we only look at a two-period scheduling problem with simultaneous decisions. Additional research is needed to study the CP game under a multi-period or a sequential context, where in the latter, agents demand shifting decisions are made stage by stage given the non-anticipatory price and demand realization. This also motivates the last limitation, which is that we did not consider the incomplete information of the game model. In practice, agents’ payoff structure should be private information that can’t be observed by others, but agents may get others’ previous decisions and system peak period to update their beliefs gradually. We would expect contextual optimization to be a promising way to infer the private payoff function from other observed features and embed it into the stochastic formulation.
Appendix Appendix A Proof of Proposition 4
Proof.
We separate three conditions to analyze the continuity and concavity due to the indicator function.
(1) We first analyze the concave game conditions. To make the game concave, the CP period can’t be changed, i.e., there must be only one period active in the entire gaming process. In this case, the indicator function is eliminated, and the payoff function returns to a quadratic function as described in (3), which is concave and continuous. Because changing CP period requires the sum of all agents’ demand shifting greater than the system balance point $b$, and the greatest demand shifting is their critical point. Thus, the conditions is $b<-r_{\mathrm{x}}-r_{\mathrm{y}}$ when $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$ and $b>r_{\mathrm{x}}+r_{\mathrm{y}}$ when $S_{\mathrm{b},1}<S_{\mathrm{b},2}$.
(2) We then prove the quasiconcave/discontinuous conditions. When the CP period changes in the gaming process, it is obvious that the indicator function makes the game discontinuous. We then write the payoff function of agent x with the switching point $c_{\mathrm{x}}=b-y$
$f_{\mathrm{x}}(x,y)$ $=-(X_{1}+x)I(x-c_{\mathrm{x}})-(X_{2}-x)I(c_{\mathrm{x}}-x)-\alpha_{\mathrm{x}}x^{2},$ (27)
According to the definition of the quasiconcave function, for all $x^{\prime},x^{\prime\prime}\in\mathcal{X}_{\mathrm{x}}$ and $\lambda\in[0,1]$, agent x’s payoff function $f_{\mathrm{x}}(x,y)$ should satisfy the following for all $y\in\mathcal{X}_{\mathrm{y}}$,
$f_{\mathrm{x}}(\lambda x^{\prime}+(1-\lambda)x^{\prime\prime},y)\geq\min\{f_{\mathrm{x}}(x^{\prime},y),f_{\mathrm{x}}(x^{\prime\prime},y)\},$ (28)
which is determined by the switching point $c_{\mathrm{x}}$ and the critical point $r_{\mathrm{x}}$. There are three cases.
i) Agent x can switch CP period before reaching critical points in both periods 1 and 2, i.e., the switching point is within the critical point $-r_{\mathrm{x}}\leq c_{\mathrm{x}}\leq r_{\mathrm{x}}$, and we have
$-r_{\mathrm{x}}\leq b-y,r_{\mathrm{x}}\geq b-y,$ (29)
Graphically, the function is monotonically increasing before the switching point and decreasing after the switching point.
ii) Agent x can only switch CP period before reaching critical points in period 1, i.e., the switching point is greater than the critical point of period 2 $r_{\mathrm{x}}<c_{\mathrm{x}}$. Also, the function is l.s.c. as defined in Definition 1, i.e., $-(X_{2}-c_{\mathrm{x}})<-(X_{1}+c_{\mathrm{x}})$, and we have
$r_{\mathrm{x}}<b-y,\ \frac{X_{2}-X_{1}}{2}=b_{\mathrm{x}}<b-y,$ (30)
Graphically, the function is quadratic (concave) in the left part of the switching point, monotonically decreasing in the right part of the switching point, and the function is l.s.c. in the switching point.
iii) Agent x can only switch CP period before reaching critical points in period 2, i.e., the switching point is less than the critical point of period 1 $-r_{\mathrm{x}}>c_{\mathrm{x}}$. Also, the function is u.s.c. as defined in Definition 1, i.e., $-(X_{2}-c_{\mathrm{x}})<-(X_{1}+c_{\mathrm{x}})$, and we have
$-r_{\mathrm{x}}>b-y,\ \frac{X_{2}-X_{1}}{2}=b_{\mathrm{x}}>b-y,$ (31)
Graphically, the function increases monotonically in the left part of the switching point, is quadratic (concave) in the right part of the switching point, and is u.s.c. in the switching point.
Combining these three scenarios, we need both agents’ payoff functions to satisfy one of them for all other agents’ strategies to make the game quasiconcave. When agent x satisfy case i), according to the condition (29), agent y’s strategy needs to satisfy $y\in[b-r_{\mathrm{x}},b+r_{\mathrm{x}}]$. From the graph description, we know agent y’s strategy is $y=c_{\mathrm{y}}=b-x$, $y=-r_{\mathrm{y}}$, and $y=r_{\mathrm{y}}$ under case i), ii), and iii), respectively. In case i), if $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}$, agent x can balance its demand in the two periods, thus the maximum $x=b_{\mathrm{x}}$, if $b_{\mathrm{x}}>r_{\mathrm{x}}$ or $b_{\mathrm{x}}<-r_{\mathrm{x}}$, the best strategy is either $r_{\mathrm{x}}$ or $-r_{\mathrm{x}}$. This means $x\in[-r_{\mathrm{x}},r_{\mathrm{x}}]$, and thus $y\in[b-r_{\mathrm{x}},b+r_{\mathrm{x}}]$. Because both agents take the best strategy at the switching point, the only solution is $x=b_{\mathrm{x}},y=b_{\mathrm{y}}$, and taking this into (29), we have the conditions
$-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},\ -r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}}.$ (32)
Otherwise, suppose agent x satisfies case ii), according to the condition (30), agent y’s strategy needs to satisfy $y<\min\{b-r_{\mathrm{x}},b-b_{\mathrm{x}}\}$. We know agent y’s strategy is $y=-r_{\mathrm{y}}$ and $y=r_{\mathrm{y}}$ under cases ii) and iii), and $r_{\mathrm{y}}$ are independent to $r_{\mathrm{x}}$ and $b_{\mathrm{x}}$. Thus, agent y under cases ii) and iii) can’t guarantee the quasiconcave conditions. We then conclude that each agent’s payoff function is quasiconcave if (32) is true, which shows each agent is capable according to Definition 3.
(3) Other than the above two conditions, the game is non-concave and still discontinuous due to the indicator function. In this case, at least one agent is non-capable, and they can change the CP period together. We can directly write the conditions as the complementary set of concave and quasiconcave conditions with respect to $\mathbb{R}$, i.e., (6d). Also, note that if one agent’s payoff function is quasiconcave or non-concave, the payoff functions of the other agents must also be quasiconcave or non-concave, which excludes the conditions that one agent’s payoff function is concave and the other is quasiconcave. This finishes the proof of this proposition. ∎
Appendix Appendix B Proof of Theorem 5
Proof.
The basic idea of proving this Theorem is to analyze whether agents are capable, if they are not, whether they are upper non-capable or lower-non-capable, and different conditions corresponding to different NE. The rationale is that both agents’ best response is to balance their demand in two periods, but this is limited by their baseline conditions, which is similar to analyzing the relationship between their critical point and balance point.
Note that the lower non-capable agent x satisfies $b_{\mathrm{x}}<-r_{\mathrm{x}}$, corresponding to scenario iii) as described in the proof of Proposition 4. Graphically, it shows the function increases monotonically in the left part of the switching point, and quadratic (concave) in the right part of the switching point, but not u.s.c. in the switching point. While the upper non-capable agent x is $b_{\mathrm{x}}>r_{\mathrm{x}}$, corresponding to scenario ii) as described in the proof of Proposition 4. Graphically, it shows the function is quadratic (concave) in the left part of the switching point, monotonically decreasing in the right part of the switching point, but not l.s.c. in the switching point.
We then separate many scenarios according to baseline demand conditions and whether they are non-capable agents, upper non-capable, or lower non-capable, to analyze the NE and corresponding conditions.
(1) $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$, i.e., baseline CP period is 1.
(1a) $X_{1}\geq X_{2},Y_{1}\geq Y_{2}$, i.e., both agents’ individual demand is higher in the baseline CP period. We have four scenarios determined by whether they are capable or not.
  • Both agents are capable agents, i.e., $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq 0,-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq 0$, which means they can balance their demand, and the NE for all agent is their balance point,
    $x^{*}=b_{\mathrm{x}},y^{*}=b_{\mathrm{y}}$ (33a)
    and we have the conditions for each agent as
    $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq 0,-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq 0.$ (33b)
  • Agent x is non-capable, while agent y is capable, i.e., $b_{\mathrm{x}}<-r_{\mathrm{x}}\leq 0,-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq 0$. Due to $X_{1}\geq X_{2}$, agent x must be lower non-capable, and the NE for agent x is the critical point, i.e., $x^{*}=-r_{\mathrm{x}}$. Agent y can at least shift demand until the system demands balance in the two periods, and as agent x can’t reach the balance point, agent y can save more by shifting more demand away from the CP period, but needs to compare with the shifting penalty. Thus, the NE for agent y is
    $y^{*}=\max\{-r_{\mathrm{y}},y^{\prime}\},y^{\prime}=b_{\mathrm{y}}-(x^{*}-b_{\mathrm{x}})=b-x^{*}\leq 0.$ (34a)
    Then we compare the $y^{\prime}$ with its critical point $-r_{\mathrm{y}}$. If $y^{\prime}<-r_{\mathrm{y}}$, i.e., $b-x^{*}=b+r_{\mathrm{x}}<-r_{\mathrm{y}}$, which is equivalent to $-r_{\mathrm{x}}-r_{\mathrm{y}}>b$, agent y can’t change the CP period solely, then $y^{*}=-r_{\mathrm{y}}$. The corresponding conditions for agent x is $-r_{\mathrm{x}}>b_{\mathrm{x}}$, and for agent y is
    $-r_{\mathrm{y}}\leq b_{\mathrm{y}}<-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}.$ (34b)
    where $-r_{\mathrm{y}}\leq b_{\mathrm{y}}$ is because agent y is capable. Otherwise, when $-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b$, the best strategy for agent y is $y^{\prime}$, and we have $y^{*}=b+r_{\mathrm{x}}$. The corresponding condition for agent x is still $-r_{\mathrm{x}}>b_{\mathrm{x}}$, for agent y is
    $b_{\mathrm{y}}\geq-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}.$ (34c)
    Note that this includes two scenarios when changing agent x to agent y.
  • Both agents are (lower) non-capable, i.e., $-r_{\mathrm{x}}>b_{\mathrm{x}},-r_{\mathrm{y}}>b_{\mathrm{y}}$, and the NE is their critical point $x^{*}=-r_{\mathrm{x}},y^{*}=-r_{\mathrm{y}}$.
(1b) $X_{1}\geq X_{2},Y_{1}<Y_{2}$, i.e., agent x’s demand is higher in the CP period, while agent y’s demand is lower. Similar to (1a), when both agents are capable, the NE is their balance point. When they are not all capable agents, their optimal strategy is determined by whether they can change the CP period together and whether they can balance their own demand. Suppose agent x is non-capable, due to $X_{1}\geq X_{2}$, agent x must be lower non-capable, i.e., $b_{\mathrm{x}}<-r_{\mathrm{x}}$. Because the CP period is 1 and the agent x’s demand is higher in period 1, although agent x is lower non-capable, it is possible to change the CP period solely without balancing its own demand. Thus, we separate two scenarios to analyze.
  • Agent x can’t change the CP period solely, i.e., $-r_{\mathrm{x}}\geq b$. Agent x’s best strategy is to shift demand away from the CP period, but can’t reach the system balance point, so agent y will also shift some demand away from the CP period to reduce its cost but must compare with the shifting penalty. Thus, the NE is
    $x^{*}=-r_{\mathrm{x}},\ y^{*}=\max\{-r_{\mathrm{y}},b-x*\}.$ (35a)
    When $-r_{\mathrm{y}}>b-x^{*}$, both agents can’t change the CP period together, and the best strategy for agent y is $y^{*}=-r_{\mathrm{y}}$. The condition for agent x is $b_{\mathrm{x}}<-r_{\mathrm{x}}$, and for agent y is $b_{\mathrm{y}}<-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}$. Otherwise, $-r_{\mathrm{y}}\leq b-x^{*}$, agent y can cooperate with agent x to change the CP period, and the best strategy for agent x is $y^{*}=b-x^{*}=b+r_{\mathrm{x}}$. The condition for agent x is the same as the lower non-capable condition $b_{\mathrm{x}}<-r_{\mathrm{x}}$, and for agent y is
    $-r_{\mathrm{x}}\geq b,\ -r_{\mathrm{x}}-r_{\mathrm{y}}\leq b,$ (35b)
    $-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}\leq b_{\mathrm{y}}\leq-r_{\mathrm{x}}-b_{\mathrm{x}}.$ (35c)
  • Agent x can change the CP period solely, i.e., $-r_{\mathrm{x}}<b$. Agent x shifts demand away from the CP period and at least changes the CP period, and agent y will move demand back to the CP period to keep the original CP period. Thus, determined by which agent can shift more before reaching their critical point limit, the NE is
    $x^{*}=\max\{-r_{\mathrm{x}},b-y^{*}\},\ y^{*}=\min\{r_{\mathrm{y}},b-x^{*}\}.$ (36a)
    When $r_{\mathrm{y}}\leq b-x^{*}$, which means agent y first reaches the critical point limit, then agent x must satisfy $b-y^{*}>-r_{\mathrm{x}}$. Thus, the best strategy is $x^{*}=b-r_{\mathrm{y}},y^{*}=r_{\mathrm{y}}$. The condition for agent x is the same as the lower non-capable condition $b_{\mathrm{x}}<-r_{\mathrm{x}}$, and for agent y is
    $b-r_{\mathrm{y}}>-r_{\mathrm{x}},\ b_{\mathrm{y}}>r_{\mathrm{y}}-r_{\mathrm{x}}-b_{\mathrm{x}}.$ (36b)
    Otherwise, we have the best strategy for both agents as $x^{*}=-r_{\mathrm{x}},y^{*}=b+r_{\mathrm{x}}$, and the corresponding condition for agent x is still $b_{\mathrm{x}}<-r_{\mathrm{x}}$, and for agent y is
    $-r_{\mathrm{x}}-b_{\mathrm{x}}<b_{\mathrm{y}}\leq r_{\mathrm{y}}-r_{\mathrm{x}}-b_{\mathrm{x}}.$ (36c)
If agent y is non-capable, due to $Y_{1}<Y_{2}$, agent y must be upper non-capable, i.e., $b_{\mathrm{y}}>r_{\mathrm{y}}$. We focus on the conditions that both agents can change the CP period together because we have shown the condition that they can’t change the CP period above. Thus, the NE is determined by which agents can shift more before reaching their critical point limit, and we have the solution structure
$y^{*}=\min\{r_{\mathrm{y}},b-x^{*}\},x^{*}=\max\{-r_{\mathrm{x}},b-y^{*}\}.$ (37a)
Following the same process with (36), we know when $-r_{\mathrm{x}}\geq b-y^{*}$, agent y must satisfy $r_{\mathrm{y}}>b-x^{*}$, and the best strategy for both agents are $x^{*}=-r_{\mathrm{x}},y^{*}=b+r_{\mathrm{x}}$. The condition for agent y is the upper non-capable condition $b_{\mathrm{y}}>r_{\mathrm{y}}$, and for agent x is
$r_{\mathrm{y}}\geq b+r_{\mathrm{x}},\ -r_{\mathrm{x}}-r_{\mathrm{y}}\leq b,$ (37b)
$-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{y}}\leq b_{\mathrm{x}}\leq-r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{y}},$ (37c)
where the second condition is because both agents can change the CP period together. Otherwise, the best strategy for both agents is $x^{*}=b-r_{\mathrm{y}},y^{*}=r_{\mathrm{y}}$, and the condition for agent y is still the upper non-capable condition $b_{\mathrm{y}}>r_{\mathrm{y}}$, and for agent x is
$-r_{\mathrm{x}}\leq b-r_{\mathrm{y}},\ r_{\mathrm{x}}+r_{\mathrm{y}}\geq b,$ (37d)
$-r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{y}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{y}}.$ (37e)
(1c) $X_{1}<X_{2},Y_{1}\geq Y_{2}$, i.e., agent y’s demand is higher in the CP period, while agent x’s demand is lower, which is the same to (1b) by changing the x to y, and we omit the redundant math here.
(2) $S_{\mathrm{b},1}<S_{\mathrm{b},2}$, i.e., baseline CP period 2. Similar to the CP period 1 condition, we can still separate three cases with (2a) $X_{1}<X_{2},Y_{1}<Y_{2}$; (2b) $X_{1}\geq X_{2},Y_{1}<Y_{2}$; and (2c) $X_{1}<X_{2},Y_{1}\geq Y_{2}$. The only difference in all cases is that when the best strategy is the critical point, the critical point changes from ±$\pm r_{\mathrm{x}}$ to $\mp r_{\mathrm{x}}$. Other analysis is the same as (1), and we omit the redundant part.
To conclude, combining all the scenarios, we obtain the NE and corresponding conditions as (7) show and prove this Theorem. ∎
Appendix Appendix C Proof of Theorem 6
Overview of the proof: The basic idea of proving this Theorem is to show that the dynamical systems of each period (8a) and (8b) are asymptotically stable in the strategy set $\mathcal{X}$, then, add the switching logic to show the system is global uniform asymptotically stable in $\mathcal{X}_{\mathrm{s}}$ as described in Theorem 6, i.e., local uniform asymptotically stable in $\mathcal{X}$. We specify the proof process as follows:
  • We first prove dynamical systems of each period (8a) and (8b) is asymptotically stable in the strategy set $\mathcal{X}$. (Lemma 16).
  • Then we prove the overall system (9) with the switching logic is global uniform asymptotically stable in $\mathcal{X}_{\mathrm{s}}$.
We first provide Lemma 16 to show the system stability within each period.
Lemma 16.
System stability of one period. The system dynamics in period 1 described by (8a) is asymptotically stable in $\mathcal{X}$, i.e., for every starting point $x,y\in\mathcal{X}$, the solution $(x(k),y(k))$ to the (8a) converges to an equilibrium point $(x^{*},y^{*})$ as $k\rightarrow\infty$, where $F_{1}(x^{*},y^{*})=0$.
Proof.
The key is to show the rate of change of $\lVert F_{1}(x,y)\rVert^{2}$ is always negative for $F_{1}(x,y)\neq 0$ [Rosen, 1965]. We have
$\frac{dF_{1}}{dk}$ $=G\frac{dx}{dk}=G\dot{x},$ (38)
where $G$ is the Jacobian of $F_{1}(x,y)$, and $G=-2\text{diag}(\alpha_{\mathrm{x}},\alpha_{\mathrm{y}})$, where $\text{diag}(\cdot):\mathbb{R}^{2}\to\mathbb{R}^{2\cdot 2}$.
Now, according to (8a) and combining with the (38), we have
$\frac{1}{2}\frac{d\lVert F_{1}^{2}\rVert}{dk}$ $=\frac{1}{2}\frac{dF_{1}^{T}F_{1}}{dk}=F_{1}^{T}\frac{dF_{1}}{dk}=F_{1}^{T}GF_{1}=\frac{1}{2}F_{1}^{T}(G+G^{T})F_{1}.$ (39a)
Because the $G+G^{T}$ is negative definite, we conclude that, for some $\epsilon>0$, (39a) is equivalent to
$\frac{1}{2}\frac{d\lVert F_{1}^{2}\rVert}{dk}\leq-\epsilon\lVert F_{1}\rVert^{2}$ (39b)
Thus, $\lim_{k\rightarrow\infty}\lVert F_{1}\rVert=0$, so that $(x(k),y(k))\rightarrow(x^{*},y^{*})$, where $(x^{*},y^{*})$ is the equilibrium point and $F_{1}(x^{*},y^{*})=0$. Following the interior trajectory theorem from [Rosen, 1965], we know $(x^{*},y^{*})\in\mathcal{X}$, which proves this Lemma by showing system (8a) is asymptotically stable.
From Lemma 16, the system (8a) is asymptotically stable in $\mathcal{X}$, and this asymptotically stable result can be extrapolated to the dynamical systems of period 2 described by (8b). We then add the switched logic to study global uniform asymptotically stability at the equilibrium point satisfied (10) in $\mathcal{X}_{\mathrm{s}}$ and prove the Theorem.
Proof of Theorem 6.
From Lemma 16, all system dynamics within each period are asymptotically stable in the strategy set $\mathcal{X}$. We then rewrite the gradient of the payoff functions in the two periods as follows,
$F_{1}=A[x,y]^{T}+C_{1},F_{2}=A[x,y]^{T}+C_{2},$ (40a)
$A=[-2\alpha_{\mathrm{x}},0;0,-2\alpha_{\mathrm{y}}],C_{1}=-[\pi,\pi]^{T},C_{2}=-C_{1}.$ (40b)
We prove the Theorem based on the multiple Lyapunov function method [Liberzon, 2003]. Since our system dynamics is liner, the basic idea is to (i) find the functions $\mathcal{V}_{j}>0$ in each system (period) $j=1,2$, for $\dot{\mathcal{V}}_{j}\neq 0$, the function $\mathcal{V}_{j}$ is always decreased along the solution of the $j$th system in the region where this system is active, i.e., it is peak; (ii) on the switching surface $S_{1}(x,y)=S_{2}(x,y)$ the function $\mathcal{V}_{j}$’s value match.
We then choose $\mathcal{V}_{j},j=1,2$ as follows:
$\mathcal{V}_{1}=-[x,y]\frac{A}{2}[x,y]^{T}-C_{1}^{T}[x,y]^{T}+d_{1},\mathcal{V}_{2}=-[x,y]\frac{A}{2}[x,y]^{T}-C_{2}^{T}[x,y]^{T}+d_{2},$ (41a)
$d_{1}=\pi S_{\mathrm{b},1},d_{2}=\pi S_{\mathrm{b},2}.$ (41b)
Note that $-\partial\mathcal{V}_{1}/\partial[x,y]^{T}=F_{1}(x,y)$ and $-\partial\mathcal{V}_{2}/\partial[x,y]^{T}=F_{2}(x,y)$.
We first show the regions that guarantees function $\mathcal{V}_{j}>0$
$\alpha_{\mathrm{x}}x^{2}+\alpha_{\mathrm{y}}y^{2}+\pi(x+y+S_{\mathrm{b},1})>0,$ (42a)
$\alpha_{\mathrm{x}}x^{2}+\alpha_{\mathrm{y}}y^{2}+\pi(-x-y+S_{\mathrm{b},2})>0,$ (42b)
which is the same as (11) shows in this Theorem. We then show the rate of change of $\dot{\mathcal{V}}_{j}$ is negative.
$\dot{\mathcal{V}}_{1}=\frac{\partial\mathcal{V}_{1}}{\partial[x,y]^{T}}F_{1}=-(\frac{A^{T}+A}{2}[x,y]^{T}+C_{1})(A[x,y]^{T}+C_{1})=[x,y]A^{\prime}[x,y]^{T}+[x,y]B^{\prime}-C_{1}^{T}C_{1},$ (43a)
$A^{\prime}=[-4\alpha_{\mathrm{x}}^{2},0;0,-4\alpha_{\mathrm{y}}^{2}],B^{\prime}=-\frac{A^{T}+A}{2}C_{1}-A^{T}C_{1}=[-4\pi\alpha_{\mathrm{x}},-4\pi\alpha_{\mathrm{y}}]^{T}.$ (43b)
We have the critical point (equilibrium point) when $\partial(\dot{\mathcal{V}}_{1})/\partial[x,y]^{T}=0$
$[x^{\prime},y^{\prime}]^{T}$ $=-(A^{\prime T}+A^{\prime})^{-1}B^{\prime}=-[\frac{\pi}{2\alpha_{\mathrm{x}}},\frac{\pi}{2\alpha_{\mathrm{y}}}]^{T}.$ (43c)
It is easy to see $A^{\prime}$ is negative definite. To show $\dot{\mathcal{V}}_{1}<0$ except the equilibrium point, we need to show $\dot{\mathcal{V}}_{1}(x^{\prime},y^{\prime})=0$, so that other point must less than zero.
$\dot{\mathcal{V}}_{1}(x^{\prime},y^{\prime})$ $=B^{\prime T}((A^{\prime T}+A^{\prime})^{-1})^{T}A^{\prime}(A^{\prime T}+A^{\prime})^{-1}B^{\prime}-B^{\prime T}((A^{\prime T}+A^{\prime})^{-1})^{T}B^{\prime}-C_{1}^{T}C_{1}$
$=-\frac{1}{4}B^{\prime T}(A^{\prime-1})^{T}B^{\prime}-C_{1}^{T}C_{1}=0.$ (43d)
In terms of $\dot{\mathcal{V}}_{2}$, we have
$\dot{\mathcal{V}}_{2}$ $=\frac{\partial\mathcal{V}_{2}}{\partial[x,y]^{T}}F_{2}=-(\frac{A^{T}+A}{2}[x,y]^{T}+C_{2})(A[x,y]^{T}+C_{2}).$ (44)
Due to $C_{2}=-C_{1}$, the structure of $\dot{\mathcal{V}}_{2}$ is the same to $\dot{\mathcal{V}}_{1}$ and we can obtain $\dot{\mathcal{V}}_{2}<0$ except the equilibrium (critical) point $[x^{\prime},y^{\prime}]^{T}=[\pi/2\alpha_{\mathrm{x}},\pi/2\alpha_{\mathrm{y}}]^{T}$, where $\dot{\mathcal{V}}_{2}(x^{\prime},y^{\prime})=0$.
Then, we show $\mathcal{V}_{1}=\mathcal{V}_{2}$ on the switching surface $S_{1}(x,y)=S_{2}(x,y)$, where $x+y=b$ and $f_{\mathrm{x},1}+f_{\mathrm{y},1}=f_{\mathrm{x},2}+f_{\mathrm{y},2}$, and the key is to show $-C_{1}^{T}[x,y]^{T}+d_{1}=-C_{2}^{T}[x,y]^{T}+d_{2}$,
$-C_{1}^{T}[x,y]^{T}+d_{1}$ $=\pi(x+y)+\pi S_{\mathrm{b},1}=\pi b+\pi S_{\mathrm{b},1}=\frac{S_{\mathrm{b},2}+S_{\mathrm{b},1}}{2},$ (45a)
$-C_{2}^{T}[x,y]^{T}+d_{2}$ $=-\pi(x+y)+\pi S_{\mathrm{b},2}=-\pi b+\pi S_{\mathrm{b},2}=\frac{S_{\mathrm{b},2}+S_{\mathrm{b},1}}{2}.$ (45b)
Then, we can conclude that $\mathcal{V}_{j}$ is always decrease except $\dot{\mathcal{V}}_{j}=0$. Also, the rate of decrease of $\mathcal{V}_{j}$ along solutions is not affected by switching, and asymptotic stability is uniform with respect to $j$.
Now, let’s analyze the convergent equilibrium point. If the game is concave, only one period could be active, i.e., period 1 is active if $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$ and vice versa. Then the function value $\mathcal{V}_{j}$ decreases over the period until $\dot{\mathcal{V}}_{j}=0$, and the system reaches the stable points. According to the Theorem 5, the equilibrium point is unique and as described in (10b) or (10c).
Otherwise, both periods will be active sequentially, and $\mathcal{V}_{j}$ will decrease over time until $\mathcal{V}_{1}=\mathcal{V}_{2}$. The reason is that the dynamics from one period always push the other period active, e.g., $F_{1}$ in period 1 always decreases $x,y$, which pushes the solution past the switching surface and active period 2. Thus, as $\mathcal{V}_{1}=\mathcal{V}_{2}$ on the switching surface, both periods are finally stable on the switching surface. This means $(x(k),y(k))\rightarrow(x^{*},y^{*})$ when $k\rightarrow\infty$, where $(x^{*},y^{*})$ is the equilibrium point. According to the Theorem 5, the equilibrium point is unique and should satisfy (10a). Thus, we prove the global uniform asymptotically stable of the overall system in (11), and the equilibrium point satisfies (10).
Appendix Appendix D Proof of Theorem 7
Proof.
The key to proving this Theorem is to select the learning rate based on the backtracking line search method. The learning rate depends on the CP periods on which the current and future steps lie, as well as the payoff functions and the gradients that each agent follows.
Suppose the current step is $h$, we express the backtracking line search condition for agent x to choose the learning rate $\tau_{\mathrm{x}}$ for the finite difference approximation (12) as follows (agent y’s learning rate $\tau_{\mathrm{y}}$ follow the same rule),
$-f_{\mathrm{x},j}(x_{h+1})$ $<-f_{\mathrm{x},j}(x_{h})-\beta_{1}\tau_{\mathrm{x},h}\lVert F_{j}(x_{h},y_{h})\rVert^{2},$ (46)
where $\beta_{1}$ is the parameter within $[0,0.5]$; $f_{\mathrm{x},j}$ means the function can take either $f_{\mathrm{x},1}$ or $f_{\mathrm{x},2}$ determined by which CP period the current and next step lies, and agent’s decision is decoupled within each period. For example, if the current step $x_{h}$ lies in period 1, $j$ will take 1 for step $h$, and if the future step $x_{h+1}$ lies in period 2, $j$ will take 2 for step $h+1$. Note that $f_{\mathrm{x}}(x)$ from (1) is formulated as a payoff (profit); here, we use $-f_{\mathrm{x}}(x)$ to express the cost.
When selecting the learning rate, we gradually reduce $\tau_{\mathrm{x},h}$ by $\beta_{2}\tau_{\mathrm{x},h},\beta_{2}\in[0,1]$ until (46) satisfy. If the next step and current step lie on the same CP period, this condition ensures the objective $-f_{\mathrm{x},j}(x_{h})$ reduces by at least $\beta_{1}\tau_{\mathrm{x},h}\lVert F_{j}(x_{h},y_{h})\rVert^{2}$. This proves the concave game convergence as all the steps lie in one period, and the objective function is concave, thus, the objective function gradually reduces until $\lVert F_{j}(x_{h},y_{h})\rVert=0$.
In terms of quasiconcave and non-concave games, there are switches during the algorithm iteration. It is easy to imagine that each agent shifts demand away from the baseline CP period at the beginning, monotonically reducing their costs, and the cost function from their baseline CP period. Once they reach the balance point (switching surface), $x_{h}+y_{h}=b$, the solution starts switching between two periods. Noted that when switching happens, both agents’ individual peak period must be different, i.e., if $X_{1}+x_{h}>X_{2}-x_{h}$ for agent x, then $Y_{1}+y_{h}<Y_{2}-y_{h}$ must hold for agent y; otherwise, there will be no switching;
We then write the difference between $-f_{\mathrm{x},1}(x)$ and $-f_{\mathrm{x},2}(x)$ as
$-f_{\mathrm{x},1}(x)-(-f_{\mathrm{x},2}(x))=\pi(X_{1}+x-(X_{2}-x)).$ (47)
Suppose $X_{1}+x_{h}>X_{2}-x_{h}$ for agent x, we know
$-f_{\mathrm{x},1}(x_{h})>-f_{\mathrm{x},2}(x_{h}),-f_{\mathrm{y},1}(y_{h})<-f_{\mathrm{y},2}(y_{h}).$ (48)
Now, consider a trajectory starting from period 1, switching to period 2, and back to period 1, i.e., $-f_{\mathrm{x},1}(x_{h}),-f_{\mathrm{x},2}(x_{h+1}),-f_{\mathrm{x},1}(x_{h+2})$, our goal is to show $-f_{\mathrm{x},1}(x)$ reduce while $-f_{\mathrm{x},2}(x)$ increase for agent x and $-f_{\mathrm{y},2}(y)$ reduce while $-f_{\mathrm{y},1}(y)$ increase for agent y through the trajectory. Due to the switching from period 1 to 2, the gradient in period 1 $F_{1}(x_{h},y_{h})$ must be negative to reduce the $(x_{h},y_{h})$ so that the CP period changes. We choose the learning rate $\tau_{\mathrm{x},h},\tau_{\mathrm{y},h}$ such that
$-f_{\mathrm{x},2}(x_{h+1})<-f_{\mathrm{x},1}(x_{h})-\beta_{1}\tau_{\mathrm{x},h}\lVert F_{1}(x_{h},y_{h})\rVert^{2},$ (49a)
$-f_{\mathrm{y},2}(y_{h+1})<-f_{\mathrm{y},1}(y_{h})-\beta_{1}\tau_{\mathrm{y},h}\lVert F_{1}(x_{h},y_{h})\rVert^{2},$ (49b)
As $X_{1}+x_{h}>X_{2}-x_{h}$ and from (48), we know (49a) is easy to be true and we use the corresponding $\tau_{\mathrm{x},h}$ to update $x_{h}$ following (46); while (49b) can’t be true, so $y_{h}$ will not be updated. Thus, the switching is caused by the update of the agent x’s decision, and we know $x_{h+1}<x_{h}$. Because the gradient is in period 1, $F_{1}<0$, we know the $x_{h+1}>-r_{\mathrm{x}}$ and suppose to be reduced to reach the critical point in period 1 until converge, which indicates the right part of the critical point in the objective function $-f_{\mathrm{x},1}$, where its gradient $-f^{\prime}_{\mathrm{x},1}>0$. Thus, we have $-f_{\mathrm{x},1}(x_{h+1})<-f_{\mathrm{x},1}(x_{h})$.
In period 2, according to the trajectory, the gradient will push the solution back to period 1, which requires $x,y$ increase, and thus, we know $F_{2}(x_{h+1},y_{h+1})>0$. We then choose the learning rate $\tau_{\mathrm{x},h+1},\tau_{\mathrm{y},h+1}$ such that
$-f_{\mathrm{x},1}(x_{h+2})<-f_{\mathrm{x},2}(x_{h+1})-\beta_{1}\tau_{\mathrm{x},h+1}\lVert F_{2}(x_{h+1},y_{h+1})\rVert^{2},$ (50a)
$-f_{\mathrm{y},1}(y_{h+2})<-f_{\mathrm{y},2}(y_{h+1})-\beta_{1}\tau_{\mathrm{y},h+1}\lVert F_{2}(x_{h+1},y_{h+1})\rVert^{2},$ (50b)
Still, from (48), we know (50a) can’t be true and (50b) is easy to realized by setting $\tau_{\mathrm{y},h+1}$. Thus, the $x_{h+1}$ will not be updated and $y_{h+1}$ will be updated and push the CP period back to 1. Following a similar analysis, we know $y_{h+2}>y_{h+1}$ and $y_{h+2}<r_{\mathrm{y}}$ and are supposed to increase to reach the critical point in period 2 until converge, which indicates the left part of the critical point in the objective function $-f_{\mathrm{y},2}$, where its gradient $-f^{\prime}_{\mathrm{y},2}<0$. Thus, we have $-f_{\mathrm{y},2}(y_{h+2})<-f_{\mathrm{y},2}(y_{h+1})$.
Now, let’s look at the entire trajectory, we have $x_{h+2}=x_{h+1}<x_{h}$ for agent x and $y_{h+2}>y_{h+1}=y_{h}$ for agent y, indicating
$-f_{\mathrm{x},1}(x_{h+2})=-f_{\mathrm{x},1}(x_{h+1})<-f_{\mathrm{x},1}(x_{h}),$ (51a)
$-f_{\mathrm{y},2}(y_{h+2})<-f_{\mathrm{y},2}(y_{h+1})=-f_{\mathrm{y},2}(y_{h}).$ (51b)
Considering trajectory starting from period 2, switching to period 1, then back to period 2 can show the $-f_{\mathrm{x},2}(x_{h+2})>-f_{\mathrm{x},2}(x_{h})$ and $-f_{\mathrm{y},1}(y_{h+2})>-f_{\mathrm{y},1}(y_{h})$ following the similar analysis, we omit the redundant math.
Now, let’s analyze if the trajectory starts from period 1 and stays more steps in period 2 before going back to period 1. Given the switching pair $\underline{h},\overline{h}$ as described in Theorem 7, i.e, $\underline{h}=\overline{h}=1,\underline{h}<h<\overline{h},h=2$. For agent y, staying in period 2 gradually increases $y$ until it goes back to period 1 or reaches the critical point of $-f_{\mathrm{y},2}$ in period 2, which means converging to the critical point. Similar to (51b), we have
$y_{\overline{h}}>y_{\overline{h}-1}>...>y_{\underline{h}+1}=y_{\underline{h}},$ (52a)
$-f_{\mathrm{y},2}(y_{\overline{h}})<-f_{\mathrm{y},2}(y_{\overline{h}-1})<...<-f_{\mathrm{y},2}(y_{\underline{h}+1})=-f_{\mathrm{y},2}(y_{\underline{h}}).$ (52b)
where the first inequality and last equality equality comes from (51b).
For agent x, although $x_{h}$ will increase in period 2, the $x_{h},y_{h}$ still within period 2, and we have
$\max\{x_{h}$ $+y_{h}|h\in(\underline{h},\overline{h})\}<\min\{x_{\overline{h}}+y_{\overline{h}},x_{\underline{h}}+y_{\underline{h}}\}$ (53a)
Because (50a) can’t be true and (50b) is easily satisfied and $y_{h}$ gradually increase in period 2 as described in (52), the system switch must be activated by agent y. This means
$y_{\underline{h}}<\max\{y_{h}|h\in(\underline{h},\overline{h})\}<y_{\overline{h}},$ (53b)
Combined with (51a), we have
$x_{\underline{h}}>\max\{x_{h}|h\in(\underline{h},\overline{h})\}=x_{\overline{h}},$ (53c)
where the first inequality is obtained due to two cases: i) if $x_{\overline{h}}+y_{\overline{h}}<x_{\underline{h}}+y_{\underline{h}}$, and we know $y_{\underline{h}}<y_{\overline{h}}$ from (53b), thus, $x_{\underline{h}}>x_{\overline{h}}$; ii) if $x_{\overline{h}}+y_{\overline{h}}\geq x_{\underline{h}}+y_{\underline{h}}$, we know (53a) is equivalent to
$\max\{x_{h}$ $+y_{h}|h\in(\underline{h},\overline{h})\}<x_{\underline{h}}+y_{\underline{h}},$ (53d)
and due to (53b), $y_{\underline{h}}<\max\{y_{h}|h\in(\underline{h},\overline{h})\}$, thus, $x_{\underline{h}}>\max\{x_{h}|h\in(\underline{h},\overline{h})\}$. Thus, according to (53c), we have
$-f_{\mathrm{x},1}(x_{\overline{h}})<-f_{\mathrm{x},1}(x_{\underline{h}})$ (53e)
This proves the Theorem by showing $-f_{\mathrm{x},1},-f_{\mathrm{y},2}$ reduce and $-f_{\mathrm{x},2},-f_{\mathrm{y},1}$ increase if $X_{1}+x>X_{2}-x$, following Theorem 6, we know the system will converge to either the switching surface $f_{\mathrm{x},1}+f_{\mathrm{y},1}=f_{\mathrm{x},2}+f_{\mathrm{y},2}$ or the critical points $-r_{\mathrm{x}}$ ($F_{1}(x,y)=0$) when $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$ or $r_{\mathrm{x}}$ ($F_{2}(x,y)=0$) when $S_{\mathrm{b},1}<S_{\mathrm{b},2}$.
Appendix Appendix E Proof of Theorem 9
Proof.
We first show the optimal solution from the centralized CP shaving model as described in (14). Under the following condition, we can easily get both agents’ solution as the critical point $r_{\mathrm{x}},r_{\mathrm{y}}$ by first-order optimality condition.
$-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b,\text{if }S_{\mathrm{b},1}\geq S_{\mathrm{b},2},$ (54a)
$r_{\mathrm{x}}+r_{\mathrm{y}}\geq b,\text{if }S_{\mathrm{b},1}<S_{\mathrm{b},2}.$ (54b)
Otherwise, the system demand will be balanced in the two periods, i.e., $x+y=b$, and we add the constraints with Lagrange multipliers $\lambda$
$\mathcal{L}(x,y,\lambda)$ $=\pi\max\{S_{1}(x,y),S_{2}(x,y)\}+\alpha_{\mathrm{x}}x^{2}+\alpha_{\mathrm{y}}y^{2}+\lambda(b-x-y),$ (54c)
$\frac{\partial\mathcal{L}}{\partial x}$ =±$=\pm\pi+2\alpha_{\mathrm{x}}x-\lambda=0,$ (54d)
$\frac{\partial\mathcal{L}}{\partial y}$ =±$=\pm\pi+2\alpha_{\mathrm{y}}y-\lambda=0,$ (54e)
$\frac{\partial\mathcal{L}}{\partial\lambda}$ $=x+y=b,$ (54f)
$x$ $=\frac{\alpha_{\mathrm{y}}}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}}b,y=\frac{\alpha_{\mathrm{x}}}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}}b,$ (54g)
where the sign of $\pm$ is determined by the system baseline demand $S_{\mathrm{b},1},S_{\mathrm{b},2}$. Overall, the solution for the centralized model (14) is
$x$ $=\frac{\alpha_{\mathrm{y}}b}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}},y=\frac{\alpha_{\mathrm{x}}b}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}},$ if $\text{if }-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}$ (55a)
$x$ $=-r_{\mathrm{x}},y=-r_{\mathrm{y}},$ if $\text{if }-r_{\mathrm{x}}-r_{\mathrm{y}}>b$ (55b)
$x$ $=r_{\mathrm{x}},y=r_{\mathrm{y}},$ if $\text{if }r_{\mathrm{x}}+r_{\mathrm{y}}<b$ (55c)
Note that the peak shaving effectiveness defined in the theorem statement is equal to directly comparing the $x^{*}+y^{*}$ with $x_{\mathrm{cen}}^{*}+y_{\mathrm{cen}}^{*}$ because the baseline demand is the same in both models. Combine with the game solution in Theorem 5, under the concave game conditions, we have $x^{*}=x_{\mathrm{cen}}^{*}=\pm r_{\mathrm{x}},y^{*}=y_{\mathrm{cen}}^{*}=\pm r_{\mathrm{y}}$, indicating the peak shaving effectiveness at equilibrium equal to 1. Under the non-concave game and quasiconcave game conditions, it is easy to see both game model and centralized model will balance system demand, i.e., $x^{*}+y^{*}=x_{\mathrm{cen}}^{*}+y_{\mathrm{cen}}^{*}=b$, meaning that the peak shaving effectiveness at equilibrium also equal to 1. This proves the theorem.
Appendix Appendix F Proof of Theorem 10
Proof.
Recall the centralized model solution from Theorem 9, the $x_{\mathrm{cen}}^{*},y_{\mathrm{cen}}^{*}$ under quasiconcave and non-concave game conditions is
$x_{\mathrm{cen}}^{*}=\frac{\alpha_{\mathrm{y}}b}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}},y_{\mathrm{cen}}^{*}=\frac{\alpha_{\mathrm{x}}b}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}}.$ (56)
Although the game solution is different under these two conditions, we can denote it as $x^{*}$, and according to the definition of PoA (18a), the PoA is
$P=\frac{f_{\mathrm{x}}(x^{*},y^{*})+f_{\mathrm{y}}(x^{*},y^{*})}{f_{\mathrm{x}}(x_{\mathrm{cen}}^{*},y_{\mathrm{cen}}^{*})+f_{\mathrm{y}}(x_{\mathrm{cen}}^{*},y_{\mathrm{cen}}^{*})}=\frac{\pi S+\alpha_{\mathrm{x}}x^{*2}+\alpha_{\mathrm{y}}y^{*2}}{\pi S+\alpha_{\mathrm{x}}(\frac{\alpha_{\mathrm{y}}b}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}})^{2}+\alpha_{\mathrm{y}}(\frac{\alpha_{\mathrm{x}}b}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}})^{2}}$ (57a)
By replacing $b=x^{*}+y^{*}$, we have
$P$ $=\frac{\pi S+\alpha_{\mathrm{x}}x^{*2}+\alpha_{\mathrm{y}}y^{*2}}{\pi S+\frac{\alpha_{\mathrm{x}}\alpha_{\mathrm{y}}^{2}(x^{*}+y^{*})^{2}+\alpha_{\mathrm{x}}^{2}\alpha_{\mathrm{y}}(x^{*}+y^{*})^{2}}{(\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}})^{2}}}=\frac{(\pi S+\alpha_{\mathrm{x}}x^{*2}+\alpha_{\mathrm{y}}y^{*2})(\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}})}{\pi S(\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}})+\alpha_{\mathrm{x}}\alpha_{\mathrm{y}}(x^{*}+y^{*})^{2}}$
$=\frac{\pi S(\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}})+\alpha_{\mathrm{x}}\alpha_{\mathrm{y}}(x^{*2}+y^{*2})+(\alpha_{\mathrm{x}}x^{*})^{2}+(\alpha_{\mathrm{y}}y^{*})^{2}}{\pi S(\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}})+\alpha_{\mathrm{x}}\alpha_{\mathrm{y}}(x^{*2}+y^{*2})+2\alpha_{\mathrm{x}}\alpha_{\mathrm{y}}x^{*}y^{*}},$ (57b)
The first two terms in the denominator and nominator are the same, thus, the difference between denominator and nominator is
$(\alpha_{\mathrm{x}}x^{*})^{2}+(\alpha_{\mathrm{y}}y^{*})^{2}-2\alpha_{\mathrm{x}}\alpha_{\mathrm{y}}x^{*}y^{*}=(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2},$ (58)
We then know $P$ will increase with $(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}$, and $P$ is continuous/differentiable function regarding $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$ and $x^{*},y^{*}$, thus $\partial P/\partial[(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}]>0$. Note that the shifting cost is $\alpha_{\mathrm{x}}x^{*2}$ for agent x and $\alpha_{\mathrm{x}}x^{*}=\partial(\alpha_{\mathrm{x}}x^{*2})/\partial x^{*}$, thus we call it marginal shifting cost. ∎
Appendix Appendix G Proof of Theorem 11
Proof.
Under the concave game condition, from Theorems 5 and 9, if $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$, we have
$x^{*}$ $=-r_{\mathrm{x}},y^{*}=-r_{\mathrm{y}},x_{\mathrm{cen}}^{*}=-r_{\mathrm{x}},y_{\mathrm{cen}}^{*}=-r_{\mathrm{y}},$ (59)
and change the critical point from $-r_{\mathrm{x}}$ to $r_{\mathrm{x}}$ will get the solution when $S_{\mathrm{b},1}<S_{\mathrm{b},2}$. Thus, $x^{*}=x_{\mathrm{cen}}^{*},y^{*}=y_{\mathrm{cen}}^{*}$ and $P=1$, indicating concave game condition is equivalent to centralized condition.
From Theorem 10, we know the PoA under quasiconcave and non-concave game conditions can be written as (57b), and nominator minus denominator is $(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}\geq 0$. Thus, $P\geq 1$ under these two conditions, indicating quasiconcave and non-concave games always cause higher (or equal) anarchy than concave games.
Then Given fixed $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$, the PoA (57b) is only affected by the solution structure $x^{*},y^{*}$. Note that $X_{1},X_{2},Y_{1},Y_{2}$ could be variant such that $X_{1}+X_{2}+Y_{1}+Y_{2}=2S$, which may cause quasiconcave or non-concave game condition, i.e., all agents satisfy the following or not,
$-\frac{\pi}{2\alpha_{\mathrm{x}}}\leq b_{\mathrm{x}}=\frac{X_{2}-X_{1}}{2}\leq\frac{\pi}{2\alpha_{\mathrm{x}}}.$ (60)
We basically fixed the other parameters that appeared in the PoA expression of (57b) to explicitly show the influence of the game type change. According to Theorem 5, we know both quasiconcave and non-concave games balance the system demand in the two period periods, i.e., $x^{*}+y^{*}=b$. For the quasiconcave game, $x^{*}=b_{\mathrm{x}},y^{*}=b_{\mathrm{y}}$, and satisfy the (60), while for non-concave game, $x^{*}=\pm r_{\mathrm{x}},y^{*}=b\mp r_{\mathrm{x}}$, and $b_{\mathrm{x}}>r_{\mathrm{x}}$ or $b_{\mathrm{x}}<-r_{\mathrm{x}}$. Suppose $x^{*}=r_{\mathrm{x}},y^{*}=b-r_{\mathrm{x}}$, then the condition is $b_{\mathrm{x}}>r_{\mathrm{x}}$ and $b_{\mathrm{y}}<b-r_{\mathrm{x}}$, indicating
$(\alpha_{\mathrm{x}}b_{\mathrm{x}}-\alpha_{\mathrm{y}}b_{\mathrm{y}})^{2}>(\alpha_{\mathrm{x}}r_{\mathrm{x}}-\alpha_{\mathrm{y}}(b-r_{\mathrm{x}}))^{2}.$ (61)
If $x^{*}=-r_{\mathrm{x}},y^{*}=b+r_{\mathrm{x}}$, the situation is similar and we omit the redundant math here. Thus, the quasiconcave game causes higher PoA than the non-concave game under fixed $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$. These finish the proof of the Theorem. ∎
Appendix Appendix H Proof of Proposition 12
Overview of the proof: We derive two lemmas to show the existence and uniqueness of NE in the multi-agent CP game $G^{\prime}$, respectively. We first show the CP game $G^{\prime}$ exists NE (Lemma 17) and show the NE exists in Lemma 17 is unique (Lemma 18). We first introduce two concepts.
(1) Payoff security (Reny [Reny, 1999]). Agent $i$ can secure the payoff $f_{i}(x_{i},x_{-i})-\epsilon\in\mathbb{R}$ at $x_{i},x_{-i}\in\mathcal{X}$ iff for every $\epsilon>0$, there exist a $\hat{x}_{i}\in\mathcal{X}_{i}$ such that $f_{i}(\hat{x}_{i},x_{-i}^{\prime})\geq f_{i}(x_{i},x_{-i})-\epsilon$ for every $x_{-i}^{\prime}$ in some neighborhood of $x_{-i}$. Furthermore, we say that a game $G^{\prime}$ is payoff secure iff every agent $i$ can secure payoff for every $x_{i}\in\mathcal{X}_{i}$.
(2) Diagonally strictly concave (Rosen [Rosen, 1965]). Define the pseudo-gradient of the sum of all agents’ payoff functions $\sum_{i\in N}f_{i}(x_{i},x_{-i})$ and total differential operator $\nabla$ as
$F(x_{i},x_{-i})=[\nabla_{1}f_{1}(x_{i},x_{-i}),\cdots,\nabla_{N}f_{N}(x_{i},x_{-i})]^{T}.$ (62a)
Then the function $\sum_{i\in N}f_{i}(x_{i},x_{-i})$ is diagonally strictly concave for $x_{i},x_{-i}\in\mathcal{X}$ if for every $x_{\mathrm{a}},x_{\mathrm{b}}\in\mathbb{R}^{N}$ and $x_{\mathrm{a}},x_{\mathrm{b}}\in\mathcal{X}$, we have
$(x_{\mathrm{a}}-x_{\mathrm{b}})F(x_{\mathrm{b}})+(x_{\mathrm{b}}-x_{\mathrm{a}})F(x_{\mathrm{a}})>0.$ (62b)
Among them, payoff security means every agent can secure a payoff value in any strategy profile if they have a strategy that provides at least this value, even if other agents slightly change their strategies. We also have the sufficient conditions for diagonally strictly concave function from Rosen [Rosen, 1965], namely that the symmetric matrix $G(x_{i},x_{-i})+G^{T}(x_{i},x_{-i})$ be negative definite for $x_{i},x_{-i}\in\mathcal{X}$, where $G(x_{i},x_{-i})$ is the Jacobian of $F(x_{i},x_{-i})$ with respect to all $x_{i}$. We then introduce the existence lemma.
Lemma 17.
Existence. The multi-agent two-period quasiconcave CP game $G^{\prime}$ as described in (19) and (20b), has a pure-strategy NE $(x_{i}^{*},x_{-i}^{*})$ as defined in (21).
Proof.
According to Reny [Reny, 1999], a compact, convex, bounded, and quasiconcave game has a pure strategy NE if the sum of the agent’s payoff functions is u.s.c. as defined in Definition 1 in the whole strategy set and the game is payoff security as defined above. Thus, we focus on (a) the sum of the agent’s payoff functions, i.e., $\sum_{i\in N}f_{i}(x_{i},x_{-i})$, is u.s.c. in $x_{i},x_{-i}\in\mathcal{X}$ and (b) the game $G^{\prime}$ is payoff secure, to prove the existence of pure-strategy NE.
(a). By summing all agent’s payoff functions,
$\sum_{i\in N}f_{i}(x_{i},x_{-i})$ $=-\pi\sum_{i\in N}(X_{i,1}+x_{i})I(S_{1}-S_{2})-\pi\sum_{i\in N}(X_{i,2}-x_{i})I(S_{2}-S_{1})-\sum_{i\in N}\alpha_{i}x_{i}^{2}$
$=\pi S_{1}I(S_{1}-S_{2})-\pi S_{2}I(S_{2}-S_{1})-\sum_{i\in N}\alpha_{i}x_{i}^{2}.$ (63)
This function is u.s.c. according to Definition 1 due to $\pi S_{1}=\pi S_{2}$ when $S_{1}=S_{2}$.
(b). We then show that each agent is payoff secure by analyzing their strategies. Given any strategy pair $(x_{i},x_{-i})$ and for any $\epsilon>0$, agents $i,-i$ can secure the payoff $f_{i}(x_{i},x_{-i})-\epsilon$ and $f_{-i}(x_{i},x_{-i})-\epsilon$. Note that here $-i$ denotes all agents but $i$.
Suppose the CP period is 1, i.e., $S_{1}\geq S_{2}$, if one of the agents among $-i$ slightly increases $x_{-i}$, the CP period is still 1, and agent $i$’s payoff is the same if agent $i$ keeps the strategy profile $x_{i}$; if one of the agents among $-i$ slightly decrease $x_{-i}$, CP period is possible to change, but agent $i$ can secure the payoff $f_{i}(x_{i},x_{-i})-\epsilon$ by slightly reducing its strategy $x_{i}$. Basically, there exist a $\delta>0$ small enough that satisfy $f_{i}(x_{i}+\delta,x_{-i}^{\prime})>f_{i}(x_{i},x_{-i})-\epsilon$ for all $S_{1}^{\prime}=X_{i,1}+x_{i}+\delta+\sum_{-i\in N}(X_{-i,1}+x_{-i}^{\prime})>S_{2}$. Thus, agents $i$ is payoff secure; a similar analysis can be used for all agents $-i$ by altering $i$, and CP period 2 also follows the same analysis. Thus, the CP game is payoff secure.
According to the Reny [Reny, 1999], the CP game has a pure-strategy NE. ∎
Lemma 18.
Uniqueness. The pure-strategy NE $(x_{i}^{*},x_{-i}^{*})$ as described in Lemma 17 is unique and obtained when (23) hold.
Proof.
According to Rosen’s method [Rosen, 1965], if the sum of the agent’s payoff function is diagonally strictly concave as defined in Definition 1, then the equilibrium point described in Lemma 17 is unique if the constraints in the strategy set are concave, i.e., $\mathcal{X}=\{x|h(x)\geq 0\}$, where $h(x)$ is a concave function.
We first study the uniqueness of pure-strategy NE in each period, with the strategy set $\mathcal{X}_{\mathrm{cp}1},\mathcal{X}_{\mathrm{cp}2}$. In strategy set $\mathcal{X}_{\mathrm{cp}1}$, we have
$\sum_{i\in N}f_{i}(x_{i},x_{-i})=-\pi\sum_{i\in N}(X_{i,1}+x_{i})-\sum_{i\in N}\alpha_{i}x_{i}^{2}.$ (64)
The Hessian $H$ with respect to $x_{i}$ is $G=G^{T}=-2\text{diag}(\alpha_{1},\cdots,\alpha_{N})$, where $\text{diag}(\cdot):\mathbb{R}^{N}\to\mathbb{R}^{N\cdot N}$. This shows that $G+G^{T}$ is obviously negative definite, thus, we know the sum of the agent’s payoff function is diagonally strictly concave. This means, according to Rosen’s method, there is a unique NE in the strategy set $\mathcal{X}_{\mathrm{cp}1}$. Following the same analysis, we can conclude there is a unique NE in the strategy set $\mathcal{X}_{\mathrm{cp}2}$.
Now, the difficult part of proving this Lemma lies in the switching between two strategy sets. Thus, we separate three cases according to the CP period to analyze the NE.
  • Case (i): CP period is always 1, i.e., $S_{1}(x_{i},x_{-i})\geq S_{2}(x_{i},x_{-i}),\forall x_{i},x_{-i}\in\mathcal{X}$;
  • Case (ii): CP period is always 2, i.e., $S_{1}(x_{i},x_{-i})<S_{2}(x_{i},x_{-i}),\forall x_{i},x_{-i}\in\mathcal{X}$.
  • Case (iii): CP period changes during the gaming process, i.e., $\exists x_{i}^{\prime},x_{-i}^{\prime},x_{i}^{\prime\prime},x_{-i}^{\prime\prime}\in\mathcal{X},S_{1}(x_{i}^{\prime},x_{-i}^{\prime})-S_{2}(x_{i}^{\prime},x_{-i}^{\prime})\geq 0,S_{1}(x_{i}^{\prime\prime},x_{-i}^{\prime\prime})-S_{2}(x_{i}^{\prime\prime},x_{-i}^{\prime\prime})<0$;
We first show Cases (i) and (ii) can’t exist under the quasiconcave conditions (20b) because only one period could be CP period during the entire solution process and the game is concave. Taking case (i) for example, mathematically, we have, $S_{1}(x_{i},x_{-i})\geq S_{2}(x_{i},x_{-i}),\forall x_{i},x_{-i}\in\mathcal{X}=\mathcal{X}_{\mathrm{cp}1}$, i.e.,
$S_{1}(x_{i},x_{-i})$ $=\sum_{i\in N}(X_{i,1}+x_{i})\geq\sum_{i\in N}(X_{i,2}-x_{i})=S_{2}(x_{i},x_{-i}),$ (65a)
$\frac{\sum_{i\in N}(X_{i,2}-X_{i,1})}{2}=b\leq\sum_{i\in N}x_{i},$ (65b)
which aligns with the conditions of the concave game. The same analysis can be applied to case (ii), where the CP period is always 2. Thus, we know both cases (i) and (ii) can’t exist under the quasiconcave conditions.
We then analyze Case (iii) and start with the two-agent setting with agent x and y. We first note that the CP charge, either $\pi(X_{1}+x)$ or $\pi(X_{2}-x)$, is relatively greater compared with penalty $\alpha_{\mathrm{x}}x^{2}$ for all agents because the CP period can change in the case (iii). From the best response perspective, suppose the CP period is 1, i.e., $S_{1}=S_{2}+\delta$, where $\delta$ is an infinitesimal number. Now, if agent x benefits by reducing $x$, which changes the CP period to 2. This indicates the demand of agent x in period 1 is greater than period 2, i.e., $X_{1}+x\geq X_{2}-x$. Now if agent y’s demand in period 1 is less than period 2, i.e., $Y_{1}+y<Y_{2}-y$, then CP period change from 1 to 2 harms agent y’s benefit, so agent y wants to push the CP period back to 1 by increasing $y$; if $Y_{1}+y\geq Y_{2}-y$, CP period changes to 2 means agent y can benefit by shifting $y$ away from period 2, i.e., still increase $y$. This means the best response of agent y is always adversarial to that of agent x.
This shows agents are fully competitive; basically, if agent x can benefit from changing strategy, that strategy harms another agent’s benefit. This is similar to the zero-sum game setting. Thus, we can apply the following min-max formulation to analyze the best responses,
$x^{*}$ $=\arg\max_{x}\min_{y}f_{\mathrm{x}}(x,y)=\arg\max_{x}\left\{\begin{aligned} -\pi(X_{1}&+x)-\alpha_{\mathrm{x}}x^{2},\\ &x\geq b_{\mathrm{x}},x\geq b-y^{*}\\ -\pi(X_{2}&-x)-\alpha_{\mathrm{x}}x^{2},\\ &x<b_{\mathrm{x}},x<b-y^{*}\end{aligned}.\right.$ (66a)
where $y^{*}$ is the optimal solution of agent y following the same structure, and the first and second case corresponds to CP periods 1 and 2. By applying the first-order optimality condition, the optimal solution of $x$ is
$x^{*}=\left\{\begin{aligned} &\max\{-r_{\mathrm{x}},b-y^{*},b_{\mathrm{x}}\},&S_{1}\geq S_{2}\\ &\min\{r_{\mathrm{x}},b-y^{*},b_{\mathrm{x}}\},&S_{1}<S_{2}\\ \end{aligned}.\right.$ (66b)
The optimal solution for agent y follows the same structure
$y^{*}=\arg\max_{y}\min_{x}f_{\mathrm{y}}(x,y)=\left\{\begin{aligned} &\max\{-r_{\mathrm{y}},b-x^{*},b_{\mathrm{y}}\},&S_{1}\geq S_{2}\\ &\min\{r_{\mathrm{y}},b-x^{*},b_{\mathrm{y}}\},&S_{1}<S_{2}\\ \end{aligned}.\right.$ (66c)
According to (5), $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}$, $-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}}$, we know if $S_{1}\geq S_{2}$, when $x^{*},y^{*}$ reach optimal solution, $y^{*}=\max\{b-x^{*},b_{\mathrm{y}}\}$ and $x^{*}=\max\{b-y^{*},b_{\mathrm{x}}\}$. Then, according to (66b) and (66c), the only solutions are $x^{*}=b_{\mathrm{x}},\ y^{*}=b_{\mathrm{y}}$, and the conditions are
$X_{2}+Y_{2}-x^{*}-y^{*}=S_{2}=X_{1}+Y_{1}+x^{*}+y^{*}=S_{1}.$ (67)
Then, we use the two-agent solution to analyze the solution in the $N$-agent setting. Given that all agents are capable, define a partition of set $N$ into two disjoint subsets, denoted as $N_{\mathrm{a}},N_{\mathrm{b}}$, such that $N_{\mathrm{a}}\cup N_{\mathrm{b}}=N,N_{\mathrm{a}}\cap N_{\mathrm{b}}=\emptyset$, and the demand for these two subsets is unequal in at least one period, i.e.,
$\{\sum_{i\in N_{\mathrm{a}}}(X_{i,1}+x_{i})\neq\sum_{i\in N_{\mathrm{b}}}(X_{i,1}+x_{i})\}\cup\{\sum_{i\in N_{\mathrm{a}}}(X_{i,2}-x_{i})\neq\sum_{i\in N_{\mathrm{b}}}(X_{i,2}-x_{i})\}.$ (68)
Then treat these two sets as two aggregated agents with strategy $x_{\mathrm{a}},x_{\mathrm{b}}$ and apply the two-agent solutions, we know the two sets will balance their demand over the two periods, i.e., the following is true for sets $N_{\mathrm{a}},N_{\mathrm{b}}$,
$x_{\mathrm{a}}^{*}=\frac{\sum_{i\in N_{\mathrm{a}}}X_{i,2}-\sum_{i\in N_{\mathrm{a}}}X_{i,1}}{2}=\frac{\sum_{i\in N_{\mathrm{a}}}(X_{i,2}-X_{i,1})}{2}=\sum_{i\in N_{\mathrm{a}}}b_{i}=\sum_{i\in N_{\mathrm{a}}}x_{i}^{*}.$ (69)
Then, applying the same partition to $N_{\mathrm{a}},N_{\mathrm{b}}$ to get the subsets $N_{\mathrm{a,a}},N_{\mathrm{a,b}}$ and $N_{\mathrm{b,a}},N_{\mathrm{b,b}}$, and applying the two-agent solutions, we still know the two subsets will balance demand over the two periods, i.e., $\sum_{i\in N_{\mathrm{a,a}}}x_{i}^{*}=\sum_{i\in N_{\mathrm{a,a}}}b_{i}$ for sets $N_{\mathrm{a,a}},N_{\mathrm{a,b}},N_{\mathrm{b,a}},N_{\mathrm{b,b}}$. Iterating applying the partition to each subset to get corresponding two subsets, the subset will finally become a singleton and we have $x_{i}^{*}=b_{i},i\in N$, and
$\sum_{i\in N}(X_{i,2}-x_{i}^{*})=S_{2}=\sum_{i\in N}(X_{i,1}+x_{i}^{*})=S_{1}.$ (70)
This shows the NE will always be obtained when $S_{1}=S_{2}$ under the quasiconcave condition, i.e., all agents’ are capable as described in Proposition 4. This also means during the game, periods 1 and 2 are active interactively, and the agent’s strategy set switches between $\mathcal{X}_{\mathrm{cp}1}$ and $\mathcal{X}_{\mathrm{cp}2}$ and finally converges at the connecting point of both strategy set $S_{1}=S_{2}$. This means there is a unique equilibrium point in the entire strategy set $\mathcal{X}$. Indeed, all agents will minimize the payment associated with $S_{1},S_{2}$ imbalance as any imbalance results in a significant CP charge change caused by the opponent’s strategy.
Thus, we conclude that the NE as described in (23) is unique and proves this Lemma. ∎
Proof of Proposition 12.
We first show the quasiconcave multi-agent CP game $G^{\prime}$ as described in (19) and (20b) has pure-strategy NE based on Lemma 17. Then, we prove the NE existed in Lemma 17 is unique based on Lemma 18 and the unique NE is obtained as described in (23). ∎
Appendix Appendix I Proof of Proposition 13
Proof.
Given $S_{\mathrm{b},1}<S_{\mathrm{b},2}$. We first assume two virtual agents as CP agent and non-CP agent. We denote the strategy as $x_{\mathrm{cp}}$, baseline demand as $X_{\mathrm{cp},2}=\sum_{i\in N_{\mathrm{cp}}}X_{i,2},X_{\mathrm{cp},1}=\sum_{i\in N_{\mathrm{cp}}}X_{i,1}$, balance point $b_{\mathrm{cp}}=(X_{\mathrm{cp},2}-X_{\mathrm{cp},1})/2$, and critical point $r_{\mathrm{cp}}=\sum_{i\in N_{\mathrm{cp}}}r_{i}$ for CP agent, and change the subscript to $\mathrm{ncp}$ for non-CP period agent. Note that for CP agent $X_{\mathrm{cp},1}<X_{\mathrm{cp},2}$ (the same definition with CP-period agent) and non-CP agent vice verse, we also have $b_{\mathrm{cp}}+b_{\mathrm{ncp}}=b$.
Then, from Theorem 5, we know these two virtual agents will balance the system demand at equilibrium. Depending on which agent reaches the critical point first, the best strategy for one agent is the critical point and the other is to balance the left unbalanced system demand. Suppose the CP agent reaches the critical point first, i.e., the demand shifting of the non-CP agent cannot be offset by the CP agent, meaning that the CP agent can’t flatten the system demand given the non-CP agent’s shifting even if the CP agent shifts the maximum amount of demand. We can express the condition as
$r_{\mathrm{cp}}<b_{\mathrm{cp}},-r_{\mathrm{ncp}}-r_{\mathrm{cp}}\leq b\leq r_{\mathrm{ncp}}+r_{\mathrm{cp}},$ (71a)
and the NE between these two virtual agents is
$x_{\mathrm{cp}}^{*}=\min\{r_{\mathrm{cp}},b_{\mathrm{cp}}\}=r_{\mathrm{cp}},x_{\mathrm{ncp}}^{*}=b-r_{\mathrm{cp}},$ (71b)
We then aggregate CP-period agents and non-CP-period agents into the CP-period agent set and non-CP-period agent set and fix the agent in each agent set, as the CP-period agent and non-CP-period agent are defined based on their baseline demand conditions. Obviously, the baseline demand, balance point, and critical point are equivalent between the agents set and the virtual agent. However, they are not totally equivalent because their strategies are different. The strategy for the CP-period agent set is determined by each agent $x_{i}^{*},i\in N_{\mathrm{cp}}$, and the non-CP-period agent set follows the same idea. Note that the same baseline demand and critical point indicate their best response rationale is the same; basically, the virtual CP agent benefits by shifting demand away from the CP period, and we have the same for the agents in the CP-period agent set.
From Theorem 5, we know each agent’s maximum shifting capacity is $b_{i}$, i.e., $x_{i}^{*}=\min\{r_{i},b_{i}\}$, and from the condition (71a), we know although the CP-period agent set and virtual CP agent have the same conditions, CP-period agent set can’t reach the best strategy of the virtual CP agent due to the following
$\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}<\sum_{i\in N_{\mathrm{cp}}}r_{i}=r_{\mathrm{cp}}=x_{\mathrm{cp}}^{*}<\sum_{i\in N_{\mathrm{cp}}}b_{i}.$ (72)
Note that the best response of both the virtual CP agent and CP-period agent set is to shift demand away from the CP period, and the benefits they obtained monotonically increase with the shifting amount. From (72), we know the maximum shifting amount of the CP-period agent set is less than the best strategy of the virtual CP agent, indicating the best strategy for the CP-period agent set is their maximum shifting amount $\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}$. In this case, the best strategy of agents in the CP-period agent set is their maximum demand shifting, i.e., $x_{i}^{*}=\min\{r_{i},b_{i}\},i\in N_{\mathrm{cp}}$
We then verify in this case the non-CP-period agent set will still balance system demand to get profits. We first show the non-concave game condition still holds, which is equivalent between virtual agents and agent set, i.e.,
$-\sum_{i\in N_{\mathrm{ncp}}}r_{i}-\sum_{i\in N_{\mathrm{cp}}}r_{i}=-\sum_{i\in N}r_{i}\leq b\leq\sum_{i\in N_{\mathrm{ncp}}}r_{i}+\sum_{i\in N_{\mathrm{cp}}}r_{i}=\sum_{i\in N}r_{i}$ (73)
Thus, under the condition (72) with $\sum_{i\in N_{\mathrm{cp}}}x_{i}^{*}=\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}$, the non-CP-period agent set can still balance system demand. Then, we show the best response rationale is the same between the non-CP-period agent set and the virtual non-CP agent due to the same baseline demand and critical point conditions. From Theorem 5, we know the virtual non-CP agent benefited by shifting demand to the CP period in response to the virtual CP agent’s strategy. Thus, the benefit of the non-CP-period agent set increases with the shifting amount to the CP period because it offsets the shifting of the CP-period agent set and further fills the system demand difference when the CP-period-agent set reaches maximum shifting capability. Also, due to $\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}<\sum_{i\in N_{\mathrm{cp}}}r_{i}$, we have $b-\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}>b-\sum_{i\in N_{\mathrm{cp}}}r_{i}$, meaning that the non-CP-period agent set can actually get more benefit than the virtual non-CP agent.
To conclude, the CP-period agents in the CP-period agent set with strategy $x_{i}^{*}=\min\{r_{i},b_{i}\},i\in N_{\mathrm{cp}}$ form a best response to the aggregated response of non-CP-period agent set with strategy $\sum_{i\in N_{\mathrm{ncp}}}x_{i}^{*}=b-\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}$ under the condition of
$\sum_{i\in N_{\mathrm{cp}}}\min\{r_{i},b_{i}\}<\sum_{i\in N_{\mathrm{cp}}}b_{i},-\sum_{i\in N}r_{i}\leq b\leq\sum_{i\in N}r_{i};$ (74)
otherwise, the best strategy will be
$x_{i}^{*}=\max\{-r_{i},b_{i}\},i\in N_{\mathrm{ncp}},\sum_{i\in N_{\mathrm{cp}}}x_{i}^{*}=b-\sum_{i\in N_{\mathrm{ncp}}}\max\{-r_{i},b_{i}\},$ (75)
corresponding to the condition
$\sum_{i\in N_{\mathrm{ncp}}}\max\{-r_{i},b_{i}\}>\sum_{i\in N_{\mathrm{ncp}}}b_{i},-\sum_{i\in N}r_{i}\leq b\leq\sum_{i\in N}r_{i}.$ (76)
Note that here, the set of agents is fixed, and we focus on the set level performance for the CP-period agent set and non-CP-period agent set. Specifically, when the aggregated (set-level) demand of the CP-period agent set is higher in the non-CP period, the CP-period agent set will perform as the non-CP-period agent set, and meanwhile, the aggregated demand of the non-CP-period agent set must be higher in the CP period, and the non-CP-period agent set perform as the CP-period agent set. This is similar to the two-agent case, where they only exchange strategy when the relative relationship between their own peak period and CP period changes.